### v1

# ClassInvGen: Class Invariant Synthesis using Large Language Models

Abstract.
Formal program specifications in the form of preconditions, postconditions, and class invariants have several benefits for the construction and maintenance of programs. They not only aid in program understanding due to their unambiguous semantics but can also be enforced dynamically (or even statically when the language supports a formal verifier). However, synthesizing high-quality specifications in an underlying programming language is limited by the expressivity of the specifications or the need to express them in a declarative manner. Prior work has demonstrated the potential of large language models (LLMs) for synthesizing high-quality method pre/postconditions for Python and Java, but does not consider class invariants.
In this work, we describe ClassInvGen, a method for co-generating executable class invariants and test inputs to produce high-quality class invariants for a mainstream language such as C++, leveraging LLMs’ ability to synthesize pure functions. We show that ClassInvGen outperforms a pure LLM-based technique to generate specifications (from code) as well as prior data-driven invariant inference techniques such as Daikon. We contribute a benchmark of standard C++ data structures along with a harness that can help measure both the correctness and completeness of generated specifications using tests and mutants. We also demonstrate its applicability to real-world code by performing a case study on several classes within a widely used and high-integrity C++ codebase.
journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
1. Introduction
Invariants are predicates that hold on the program state for all executions of the program. Many invariants hold only at specific code locations. For sequential imperative programs, it is useful to associate invariants with entry to a method (preconditions), exit from a method (postconditions), and loop headers (loop invariants). Further, for stateful classes, class invariants are facts that hold as both preconditions and postconditions of the public methods of the class, in addition to serving as a postcondition for the class constructors for the class.
These program invariants help make explicit the assumptions on the rest of the code helping modular review, reasoning, and analysis. Program invariants are useful for several aspects of software construction and maintenance during the lifetime of a program. First, executable program invariants can be enforced at runtime, thus serving as an early indicator of state corruption helping root causing and stopping a program instead of producing unexpected values. Runtime invariants serve as additional test oracles to amplify testing efforts to catch subtle bugs related to state corruption; this in turn helps with regression testing as the program evolves to satisfy new requirements. The utility of program invariants has led to design-by-contract in languages such as Eiffel (Meyer, 1992), as well as support in other languages such as Java (JML (Burdy et al., 2005)) and .NET (Code Contracts (Fähndrich, 2010)). Further, for languages that support static formal verification (e.g., Dafny (Leino, 2010), Verus (Lattuada et al., 2023), F* (Swamy et al., 2011), Frama-C (Kirchner et al., 2015)), these invariants can serve as the specification as well as help make formal verification scale by making verification modular and scalable. Unfortunately, invariants are underutilized because they require additional work and are sometimes difficult to write, so it would be useful to find a way to generate them automatically.
We focus specifically on automating the creation of class invariants for mainstream languages without first-class specification language support (e.g., C++) for several reasons:
  • Class invariants are crucial to maintain the integrity of data structures and help point to state corruption that may manifest much later within the class or in the clients. Documenting such implicit contracts can greatly aid the understanding for maintainers of the class.
  • Class invariants often form important parts of preconditions and postconditions for high-integrity data structures. Capturing such invariants in a method and asserting them in preconditions and postconditions help reduce bloat in the specifications.
  • Class invariants are challenging for users to write as writing them requires global reasoning across all the public methods for the class.
For example, consider the class for a doubly linked list as implemented in Z3 Satisfiability Modulo Theories (SMT) code [^1] https://github.com/Z3Prover/z3/blob/master/src/util/dlist.h.
1 void insert_before(T* other) {
2 ...
3 SASSERT(invariant());
4 SASSERT(other->invariant());
5 ...
6 T* prev = this->m_prev;
7 T* other_end = other->m_prev;
8 prev->m_next = other;
9 other->m_prev = prev;
10 other_end->m_next = static_cast<T*>(this);
11 this->m_prev = other_end;
12 ...
13 SASSERT(invariant());
14 SASSERT(other->invariant());
15 ...
16 }
17 bool invariant() const {
18 auto* e = this;
19 do {
20 if (e->m_next->m_prev != e)
21 return false;
22 e = e->m_next;
23 }
24 while (e != this);
25 return true;
26 }
Figure 1. An invariant in the doubly linked list class in Z3.
We see that the invariant is repeated four times: as a precondition and postcondition for the object instance this and for other. The invariant is also non-trivial, requiring local variables and a loop.
Synthesizing program invariants has been an active line of research, with both static and dynamic analysis-based approaches. Static analysis approaches based on variants of abstract interpretation (Cousot and Cousot, 1977) and interpolation (Henzinger et al., 2004) create invariants that are sound by construction. However, such techniques do not readily apply to mainstream programming languages with complex language constructs or require highly specialized methods that do not scale to large modules, since the invariants need to be additionally provably inductive to be retained. On the other hand, Daikon (Ernst et al., 2007) and successors learn invariants dynamically by instantiating a set of templates and retaining the predicates that hold on concrete test cases. While applicable to any mainstream language, it is well known that Daikon generated invariants overfit the test cases and are not sound for all test cases (Polikarpova et al., 2009). Recent works have studied fine-tuning large language models (LLMs) to learn program invariants (Pei et al., 2023) but these methods inherit the same limitation from Daikon because their training data consists of Daikon generated invariants. More importantly, the approach has not been evaluated on stateful classes to construct class invariants.
Recent work on prompting LLMs such as GPT-4 to generate program invariants for mainstream languages (Endres et al., 2024; Greiner et al., 2024; Ma et al., 2024) has been used to generate preconditions, postconditions, and loop invariants, but these methods do not readily extend to generating class invariants. Further, these methods cannot construct expressive invariants that require iterating over complex linked data structures (such as in Figure 1) other than simple arrays.
In this work, we introduce ClassInvGen, a novel method for generating high-quality object invariants for C++ classes through co-generation of invariants and test inputs using LLMs such as GPT-4o. We leverage LLMs’ ability to generate code to construct invariants that can express properties over complex data structures. The ability to consume not only the code of a class but also the surrounding comments and variable names helps establish relationships difficult for purely symbolic methods. Since an LLM can generate incorrect invariants, the method also generates test inputs to heuristically prune incorrect candidate invariants.
We leverage the framework proposed by Endres et al. (Endres et al., 2024) to evaluate the test-set correctness and completeness given a set of hidden validation tests and mutants. We contribute a new benchmark comprising standard C++ data structures along with a harness that can help measure both the correctness and completeness of generated invariants (Section 4.1). We demonstrate that ClassInvGen outperforms a pure LLM-based technique for generating program invariants from code (Sections 4.34.4) as well as prior data-driven invariant inference techniques such as Daikon (Section 4.5). We also demonstrate its applicability for real-world code by performing a case study on ema, dlist, heap, hashtable, permutation, scoped_vector, and the most complex class, bdd_manager within the Z3 SMT solver codebase; the developers of the codebase confirmed most of the new invariants proposed by ClassInvGen for these modules (Section 5).
Our contributions are summarized below:
  • Present new techniques of invariant-test co-generation by combining simple static analysis with LLM and implement an end-to-end prototype (Section 3).
  • Introduce a high-quality ClassInvGen-instrumented benchmark for evaluating object invariants (Section 4.1).
  • Investigate LLM-assisted class invariant synthesis (Sections 4.34.5).
  • Conduct a case study on Z3 bdd_manager module using ClassInvGen (Section 5).
2. Running Example: AVL Tree
Throughout this paper, AvlTree (Wikipedia contributors, 2024; Dimitrios, 2023) from our benchmark (Section 4.1) will be used as a running example to illustrate the workflow. An AVL tree is a self-balancing binary search tree (BST) where the difference in heights between the left and right subtrees of any node (called the balance factor) is at most 1. This ensures that the tree remains approximately balanced. In this implementation shown in Figure 2, the AvlTree class contains several public methods: insert, remove, contains, clear, height, size, empty, in_order_traversal, pre_order_traversal, post_order_traversal. AvlTree maintains several class invariants determined by the authors of this paper:
  • BST Property: Left child values are less than the node’s value, and right child values are greater.
  • Balance Factor: For each node, the difference in the heights of the left and right subtrees is between -1 and 1.
  • Correct Heights: The height of each node is 1 plus the maximum height of its children.
Each of these invariants should hold true before and after every public method call, and be established after the constructor method. The task is to infer these high-quality invariants from the source code.
1class AvlTree {
2public:
3 AvlTree();
4 AvlTree(const AvlTree &t);
5 AvlTree &operator=(const AvlTree &t);
6 ~AvlTree();
7
8 void insert(const T &v);
9 void remove(const T &v);
10 bool contains(const T &v);
11 void clear();
12 int height();
13 int size();
14 bool empty();
15 std::vector<T> in_order_traversal() const;
16 std::vector<T> pre_order_traversal() const;
17 std::vector<T> post_order_traversal() const;
18private:
19 struct Node {
20 T data;
21 std::unique_ptr<Node> left;
22 std::unique_ptr<Node> right;
23 int balance_factor();
24 };
25// rest of the file
26}
Figure 2. Header file of AvlTree.
3. Approach
Figure 3. Overview of ClassInvGen.
An overview of the ClassInvGen framework is shown in Figure 3. It outlines an automated pipeline for inferring class invariants from source code. ClassInvGen takes a complete source program as input and outputs invariant candidates it has identified with high confidence (called filtered invariants). ClassInvGen starts with a preprocessing step which performs static analysis on the program (Section 3.1.1). Next, an LLM is used to generate candidate invariants and filtering test suites (Section 3.1.2). Then, the code is instrumented to facilitate checking candidate invariants (Section 3.1.3). Finally, ClassInvGen uses generated tests to prune invariant candidates (Section 3.2), and a refinement loop is used to iteratively improve the results (Section 3.3).
3.1. Generation
3.1.1. Preprocessing
As illustrated in Figure 3, the generation phase begins with a static analysis of the source program. ClassInvGen uses a Tree-Sitter-based parser for program preprocessing; Tree-Sitter (Brunsfeld et al., 2024) is a parser generator tool that constructs a syntax tree from source files.
ClassInvGen parses the entire source program into an abstract syntax tree (AST) to extract class members and their recursive dependencies. It then identifies the target class and gathers details (e.g., method declarations, field declarations, and subclass definitions) relevant to forming correct class invariants. ClassInvGen recursively analyzes all identified classes (i.e., the target class and its subclasses) and performs a topological sort to prepare generation from the leaf class upward as shown in Algorithm 1.
1
2
3
4Function GenerateInvariant(target_class):
5       if target_class.id $\in$ invariants_dict then
6             return invariants_dict[target_class.id]
7
8
9      dep_classes $\leftarrow$ getClassRecursively(target_class)
10       rev_topsorted_classes $\leftarrow$ reverseToplogicalSort(dep_classes)
11
12      foreach class $\in$ rev_topsorted_classes do
13
14            class_code $\leftarrow$ getCodeForClass(class, invariants_dict)
15
16            foreach dep $\in$ getClassDependencies(class) do
17                   // Invariant: dep has been generated invariants for
18                   dep_code $\leftarrow$ getCodeForClass(dep, invariants_dict)
19                   class_code $\leftarrow$ class_code + dep_code
20
21
22            invariants_dict[class.id] $\leftarrow$ generateInvariantWithLLM(class_code)
23
24
25      final_invariants $\leftarrow$ invariants_dict[target_class.id]
26       return final_invariants
27
28
29Function getCodeForClass(class, invariants_dict, include_method_bodies=False):
30       class_text $\leftarrow$ class.get_declaration_text()
31        // Header code
32       if class.id $\in$ invariants_dict and invariants_dict[class.id] then
33             class_text $\leftarrow$ invariants_dict[class.id] + class_text
34              // Get generated invariants
35
36
37      if include_method_bodies then
38             // Add methods bodies to class_text if context allows
39
40      return class_text
41
Algorithm 1 Function to generate invariants for target class AST parsed by Tree-Sitter
3.1.2. Generation by LLM
1You are an expert in creating program invariants from code and natural language.
2Invariants are assertions on the variables in scope that hold true at different program points
3We are interested in finding invariants that hold at both start and end of a function within a data structure. Such an invariant is commonly known as an object invariant.
4
5The invariants can usually be expressed as a check on the state at the particular program point. The check should be expressed as a check in the same underlying programming language which evaluates to true or false. To express these, you can use:
6- An assertion in the programming language
7- A pure method (which does not have any side effect on the variables in scope) that checks one or more assertion
8- For a collection, you can use a loop to iterate over elements of the collection and assert something on each element or a pair of elements.
9
10Task Description:
11Task 1: Given a module, in the form of a class definition, your task is to infer object invariants about the class. For doing so, you may examine how the methods of the class read and modify the various fields of the class.
12For coming up with invariants, you may use the provided code and any comments in the code. You may also use world knowledge to guide the search for invariants.
13
14Task 2: Generate unit tests for the class based on the class definition and public API methods. The test cases should simulate a series of public method calls to verify the behavior of the class, but do not use any testing framework like gtest. Do not add assert or any form of assertions.
Figure 4. ClassInvGen Generation system prompt: instruction and task description.
After building the source program AST, ClassInvGen uses LLMs to analyze the class module and infers both invariants for the target class and tests that exercise the class’s implementation as thoroughly as possible. ClassInvGen uses a fixed system prompt that defines class invariants and outlines two main tasks: (1) generating class invariants from the source code, and (2) creating a test suite of valid API calls without specifying expected outputs (Figure 4 and Figure 23 in Appendix). Next, ClassInvGen instantiates a user prompt template (Figure 24 in Appendix) with the actual target class.
From the source program AST, ClassInvGen identifies program dependencies and populates the prompt template with the leaf struct/class. Starting from the leaf nodes, ClassInvGen leverages previously generated invariants by including them in the prompt for later classes. To accommodate the LLM’s context window limit, only the relevant child classes of the current target class are included in the prompt, with method implementations and private fields/methods hidden when necessary. An algorithm for this process is presented in Algorithm 1.
Algorithm 1 presents the invariant generation process for a source program AST. The main function GenerateInvariant takes a target_class and leverages a caching mechanism through invariants_dict to avoid redundant computations (Line 1). The algorithm first collects dependent classes via getClassRecursively (Line 1) and sorts them using reverseToplogicalSort to ensure dependency-aware processing (Line 1). For each class in the sorted order, it constructs the necessary context by obtaining the class code through getCodeForClass (Line 1). For each dependency dep of the current class, it retrieves the dep_code and concatenates it with class_code (Lines 11). The algorithm then generates invariants using generateInvariantWithLLM and stores them in invariants_dict (Line 1). The helper function getCodeForClass constructs class representations by combining the declaration text with any existing invariants from invariants_dict (Lines 11). It optionally includes method bodies based on context window constraints. This approach ensures efficient invariant generation while maintaining all necessary context and dependencies (Line 1). The algorithm concludes by returning final_invariants for the target class, effectively managing the invariant generation process while respecting LLM context limitations.
ClassInvGen accommodates large codebases by dividing the source program into smaller modules that fit within the LLM’s context window. It then iteratively generates invariants and test cases, starting from leaf classes and working up towards the root class. At each step, ClassInvGen leverages previous invariants generated for child classes to inform the invariants for parent classes.
For the AvlTree example, we begin by instantiating the prompt with Node for annotation, followed by AvlTree, since Node is a subclass of AvlTree, as illustrated in Figure 2. In this specific case, however, Algorithm 1 does not make a difference due to the small size of the source program; the entire AvlTree code fits within the LLM’s context window easily. In contrast, when working with the bdd_manager class in Z3 (around 1700 lines of code), ClassInvGen begins generation with bdd, a subclass of bdd_manager. Algorithm 1 enables ClassInvGen to partition bdd_manager class and outputs meaningful class invariants (see Section 5).
3.1.3. Instrumentation
1bool AvlTree::empty() {
2 check_invariant();
3 auto ret = empty_original();
4 check_invariant();
5 return ret;
6}
7
8bool AvlTree::empty_original() { return n == 0; }
Figure 5. AvlTree Instrumented with invariants
1void AvlTree::check_invariant() {
2 std::function<bool(const std::unique_ptr<Node>&)> is_balanced = [&](const std::unique_ptr<Node>& node) -> bool {
3 if (!node) return true;
4 int left_height = height(node->left);
5 int right_height = height(node->right);
6 if (std::abs(left_height - right_height) > 1) return false;
7 return is_balanced(node->left) && is_balanced(node->right);
8 };
9 assert(is_balanced(root));
10
11}
Figure 6. Example of a correct AvlTree class invariant
1void AvlTree::check_invariant() {
2 assert(height(root) == get_height(root));
3}
Figure 7. Example of an incorrect invariant of AvlTree because there is no get_height method
To check candidate invariants, each public method is automatically instrumented with a check_invariant call at both the start and end of its implementation, as shown in Figure 5. ClassInvGen instruments these invariants to use the compiler to rule out syntactic/type-based errors in the invariants. Each invariant is implemented as a method call to prevent conflicts with local variables within the two invocations to check_invariant. To check a specific invariant, the code for the invariant is plugged in as the implementation for the check_invariant function, and assertions are added to check that the invariant holds, as depicted in Figures 6 and 7. This setup ensures that during pruning, described below, whenever a public API call is made, each invariant candidate is automatically checked.
3.2. Heuristic Pruning
As discussed above, the LLM generates test suites for candidates. ClassInvGen wishes to select the suite that is most effective at identifying inaccurate invariants. To select the most suitable test suite, we employ line coverage as a metric. Line coverage is used to select the best test suite because it serves as a straightforward proxy for test suite completeness. The tests with the highest coverage are considered filtering tests and are used to prune the candidate invariants.
Notably, when generating tests, ClassInvGen focuses on creating valid sequences of API calls to exercise the class under test but does not assert the expected behavior. This approach simplifies the process by eliminating the need for a test oracle, which is unnecessary in the ClassInvGen framework since the purpose of test generation is to filter invariant candidates, not to test the source program directly.
[…]
These filtering tests, which are assumed to contain valid sequences, are then used to identify high-quality invariant candidates, referred to as filtered invariants. If an invariant candidate successfully compiles and runs with the filtering tests, it is designated a filtered invariant and the final output of ClassInvGen.
Figure 6 illustrates a correct filtered invariant, stating that the height of an AvlTree is balanced. In contrast, Figure 7 shows an incorrect invariant that was filtered out by the filtering tests during compilation because it references a get_height method, which is not accessible to the user.
To filter invariants, ClassInvGen compiles and runs candidate invariants with the filtering tests, retaining those that do not cause crashes and adding them to the set of filtered invariants.
1int main() {
2 // Test Case 1
3 // Basic insertions and traversals
4 {
5 AvlTree tree;
6 tree.insert(10);
7 tree.insert(20);
8 tree.insert(5);
9 tree.in_order_traversal();
10 tree.pre_order_traversal();
11 tree.post_order_traversal();
12 }
13 // Test Case 2
14 // Checking size, height, and empty status
15 {
16 AvlTree tree;
17 tree.insert(10);
18 tree.insert(20);
19 tree.insert(5);
20 tree.size();
21 tree.height();
22 tree.empty();
23 }
24 ... // the rest tests are skipped
25}
Figure 8. A test suite generated for AvlTree
3.3. Refinement
Invariants that fail at either compilation or runtime are candidates for refinement loops. For such failing invariants, ClassInvGen implements a feedback loop by collecting compiler output, error messages, and the results of filtering tests, then feeding this information back to the LLM using a dedicated prompt template (Figures 2526 in Appendix) to generate refined invariants. The system prompt for refinement is fixed, while the user prompt is instantiated with the failing invariant, compiler error messages, the target class’s source code, and the filtering tests that caused the failure.
For instance, the GCC error output shown in Figure 9 for the AvlTree candidate invariant in Figure 10 leads to the refined candidate invariant in Figure 11. We set the default threshold for refinement loops to 3, balancing the cost of calling LLMs with the benefit of repairing a single invariant; empirically, this number works well, though it may require adjustment for different codebases. If the invariant still fails after 3 refinement attempts, ClassInvGen discards it and moves on. Invariants are accepted as filtered only after successfully passing all filtered tests.
1avl_tree.cpp: In lambda function:
2avl_tree.cpp:21:16: error: use of ’is_balanced’ before deduction of ’auto’
Figure 9. gcc compiler error messages
1void AvlTree::check_invariant() {
2  auto is_bst = [&](const std::unique_ptr<Node>& node, const T& min, const T& max) -> bool {
3 if (!node) return true;
4 if (node->data <= min $|$ node->data >= max) return false;
5 return is_bst(node->left, min, node->data) && is_bst(node->right, node->data, max);
6 };
7
8 assert(is_bst(root, std::numeric_limits<T>::min(), std::numeric_limits<T>::max()));
9}
Figure 10. BST property before refinement: the word auto is causing the compiler to complain
1void AvlTree::check_invariant() {
2  std::function<bool(const std::unique_ptr<Node>&)> is_bst = [&](const std::unique_ptr<Node>& node, const T& min, const T& max) -> bool {
3 if (!node) return true;
4 if (node->data <= min $|$ node->data >= max) return false;
5 return is_bst(node->left, min, node->data) && is_bst(node->right, node->data, max);
6 };
7
8 assert(is_bst(root, std::numeric_limits<T>::min(), std::numeric_limits<T>::max()));
9}
Figure 11. BST property after refinement: auto is changed to explicit declarations. This change resolved type ambiguity, allowing the code to compile successfully.
4. Results
4.1. Benchmark
Table 1. Characteristics of the benchmark data structures. # LoC represents lines of code, # methods indicates the number of implemented methods, and # dep. shows the number of dependent classes for each data structure.
avl_ binary_ hash_ heap linked_ queue red_black_ stack vector
tree search_tree table list tree
# LoC 249 229 176 135 172 117 282 96 117
# methods 25 22 11 14 14 12 25 11 18
# dep. 1 1 0 0 0 0 1 0 0
We use a C++ implementation of classical data structures for our micro-benchmarks (Dimitrios, 2023), which include 9 data structures with associated unit tests: AvlTree, BinarySearchTree, HashTable, Heap, LinkedList, Queue, RedBlackTree, Stack, and Vector. Table 1 shows the statistics of these benchmark data structures. To ensure correctness, we thoroughly examined each benchmark example and corrected a few implementation bugs, treating this refined benchmark as the ground truth. All subsequent experiments are based on this benchmark setup.
In addition to textbook examples, we also conducted experiments on utility classes (z3Prover, 2024) from Z3 (de Moura and Bjørner, 2008), including ema, dlist, heap, hashtable, permutation, scoped_vector, and the most complex class, bdd_manager. The latter will be discussed in detail as a case study, including an evaluation with one of the authors of Z3.
4.2. Evaluation
In this section, we evaluate the quality of ClassInvGen invariants. Specifically, we explore the following research questions:
  1. (1)
    How many of these invariants are correct (with respect to the user-provided test cases) and do they capture essential properties of the source code (Section 4.3)?
  2. (2)
    How complete are the invariants in their ability to distinguish the correct program from buggy counterparts (Section 4.4)?
  3. (3)
    How does ClassInvGen compare to a state-of-the-art technique in invariant generation (namely Daikon, the most widely adopted tool for dynamic invariant synthesis) (Section 4.5)?
Our experiments were conducted on a machine with 24 CPU cores and 64 GB of RAM. We implemented ClassInvGen using GPT-4o as the underlying LLM, with its default temperature setting of 1.
Figure 12. Evaluation of ClassInvGen generated invariants
4.3. Correctness
ClassInvGen produces filtered invariants, which we evaluate using an automated pipeline (Figure 12) to assess their quality against the data structures benchmark described in Section 4.1. The evaluation process begins by instrumenting the source code with the candidate invariants, as detailed in Section 3.1.3. Unlike the generated tests used in ClassInvGen for filtering, the evaluation relies on unit tests containing input-output pairs. A filtered invariant is considered correct if it does not report any errors for any tests that successfully compile and run. Notably, these unit tests are not available during the ClassInvGen generation phase and are used exclusively for evaluation purposes.
For each passing filtered invariant, the authors of the paper conducted a manual review. This examination confirmed that all validated filtered invariants are indeed correct and effectively capture properties of the data structures. e.g., all 3 properties of AvlTree described in Section 2 are found in ClassInvGen invariants.
[…]
# invariants # compiles # pass unit tests unit test pass rate
avl_tree 30 28 17 56.7%
queue 30 30 30 100.0%
linked_list 32 32 24 75.0%
binary_search_tree 25 25 24 96.0%
hash_table 29 29 26 89.7%
heap 22 22 12 54.5%
red_black_tree 26 26 15 57.7%
stack 15 15 11 73.3%
vector 18 18 16 88.9%
Average 25.22 24.56 19.44 77.0%
[…]
Example # filtering tests # filtered inv. # good inv. unit tests pass rate
(coverage) (1 refine) (1 refine)
avl_tree 10 (91.7%) 15 20 100.0%
queue 8 (100.0%) 20 29 100.0%
linked_list 8 (92.3%) 22 36 100.0%
binary_search_tree 12 (95.6%) 16 19 100.0%
hash_table 9 (90.3%) 25 27 100.0%
heap 7 (96.0%) 10 11 100.0%
red_black_tree 13 (85.4%) 11 18 100.0%
stack 8 (100.0%) 14 14 100.0%
vector 10 (94.6%) 22 22 100.0%
Average 9.44 (94.0%) 17.22 21.78 100.0%
Table 3. For each example, the table shows the number and line coverage of filtering tests, the number of filtered invariants without refinement, and with 1 refinement, as well as the unit test pass rate after 1 refinement. With filtering tests and 1 refinement, ClassInvGen achieves a perfect unit test pass rate.
4.4. Completeness
Table 4. ClassInvGen Performance Over Baseline for Previously Survived Mutants. The table shows additional mutants killed by ClassInvGen compared to the baseline and percentage improvement.
Data Structure Unsolved Baseline (#(%)) Add. Solved by ClassInvGen (#) Improv. (%)
binary_search_tree 107(23.67) 7 6.54
hash_table 258(37.83) 38 14.73
heap 108(32.24) 12 11.11
linked_list 57(13.54) 2 3.51
red_black_tree 184(27.10) 9 4.89
stack 67(28.39) 6 8.96
vector 101(29.79) 33 32.67
avl_tree 84(17.57) 0 0.00
queue 91(26.30) 11 12.09
Total 1057 118 11.16
To evaluate the completeness of ClassInvGen invariants, we use a variation of mutation testing used in prior work to evaluate the strength of correct specifications (Endres et al., 2024). Mutation testing creates copies of the code with small changes. Individual copies are called mutants. If a mutant fails tests, the mutant is said to be killed, otherwise it survives. We expect that many mutations will corrupt data structures, so the number of mutants killed by ClassInvGen invariants is an indication of the strength of the invariant, or how complete those invariants are. Note that we only consider invariants that are correct.
We use the freeware tool mutate_cpp (Lohmann and other contributors, 2017) to perform program mutation, producing between 236 and 682 mutants per program. There are three possible outcomes for each mutant: it fails to compile, it crashes during execution, or it compiles and runs successfully without errors. We focus exclusively on mutants that either compile successfully but crash during execution or survive execution without error, as these are distinguishable cases. Mutants that fail to compile remain consistent across all configurations and therefore do not contribute to meaningful differentiation.
Figure 13. Completeness Experiment Result. The 3 bars from left to right are Tests, ClassInvGen, Tests+ClassInvGen. Tests+ClassInvGen kills the most mutants.
We conducted experiments to evaluate the completeness of three configurations: unit tests only, ClassInvGen only, unit tests with ClassInvGen. The last configuration has the strongest test oracles. For ClassInvGen only, we removed all assertions checking the expected output in the unit tests, leaving only the sequence of method calls consistent with the original tests.
[…]
Figure 13 indicates tests with ClassInvGen kill the most mutants, showing that instrumenting unit tests with ClassInvGen invariants enhances their ability to identify buggy implementations. Figure 14 shows HashTable a mutant that survived tests but killed by ClassInvGen invariant assert(this->_size > 0 || this->table.empty()). Figure 15 shows another example killed by ClassInvGen invariant assert(load_factor > 0).
1 void HashTable::clear_table() {
2 this->table.clear();
3 this->_num_elements = 0;
4 - this->_size = 0;
5 + this->_size += 0;
6 }
7}
Figure 14. Mutant that survived unit tests but killed by ClassInvGen
1 this->hash_function = hash_function;
2 this->_num_elements = 0;
3 this->_size = size;
4 - this->load_factor = 0.75;
5 + this->load_factor = -0.75;
6 this->table =
7 std::vector<std::shared_ptr<std::vector<std::pair<Key, Value>>>>(size);
8 }
Figure 15. Another Mutant that survived unit tests but killed by ClassInvGen
However, mutation testing has its limitations. For example, a mutant may be equivalent to the original program (e.g., n = 0 and n &= 0), resulting in no observable behavioral difference. Additionally, certain mutants may fall outside the scope of the invariants to detect, such as when only observer methods are altered, which do not affect the program’s state directly.
4.5. Comparison of ClassInvGen v.s. Daikon
We compare ClassInvGen with Daikon, the most widely adopted dynamic invariant generator (Ernst et al., 2007).
We replaced the invariant synthesis component in Figure 3 with Daikon, using filtering tests to generate program traces. We then ran Daikon’s invariant detector on these traces with a default confidence limit equal to 0.99 (larger values yield stronger filtering). For the most comprehensive set of invariant candidates, we combined all invariants Daikon generated at each function’s ENTER and EXIT points and used ClassInvGen invariant instrumentation to identify filtered invariants using unit tests to filter. The final output is a set of Daikon invariants that successfully pass unit test execution. On average, each benchmark example has around 5 Daikon invariants; however, in some cases, 0 to 2 of these invariants may be incorrect as shown in Table 6 examined by humans.
It is important to note that Daikon invariants are not always correct, even with respect to the tests used to infer them. Daikon outputs an invariant if its confidence level exceeds a specified configuration threshold, regardless of whether the invariant holds universally. Through the manual review, we identified 7 incorrect class invariants, as summarized in Table 5. These invariants pass unit test validation because both the filtering tests and the unit tests coincidentally constructed data structures with similar parameters , producing similar program traces. Consequently, validation was unsuccessful because these Daikon invariants based on filtering test traces also happened to hold true on the unit tests.
For instance, when Vector objects are constructed in both filtering tests and ground truth unit tests, elements are consistently push_backed in increasing order. However, users of Vector are not restricted to this behavior and could, in practice, choose to push_back larger elements first, breaking this assumed invariant. Therefore, this->data[] sorted by < is not a true class invariant.
Table 5. Incorrect Daikon Invariants that pass unit tests validation.
Data Structure Invariant Description
HashTable this->_size one of {5, 10, 40}
Heap this->data._Vector_base< int, std::allocator<int> >. _M_impl._Vector_impl_data._M_start[] elements >= 1
Vector this->data[] sorted by <
Queue this->maxSize == 10
LinkedList this->tail[].data elements >= 1
Stack this->maxSize one of {10, 160, 1280}
this->maxSize == 10
Table 6. Daikon Incorrect Invariants per Benchmark
Data Structure Total # Invariants Incorrect Invariants
hash_table 8 1
binary_search_tree 3 0
heap 10 1
red_black_tree 2 0
avl_tree 4 0
vector 3 1
stack 6 2
queue 7 1
linked_list 4 1
Average 5.2 0.78
Figure 16. Daikon vs. ClassInvGen Kills
Among the 40 correct class invariants generated by Daikon and reviewed by us, the most common invariants indicate that the class pointer is not null and remains unchanged throughout the object’s lifetime (27 invariants). The second most frequent type asserts that the number of elements (typically this->n) is non-negative (6 invariants). These two categories provide minimal utility in distinguishing and eliminating mutants (incorrect programs).
The most valuable invariants are those that capture meaningful constraints, such as this->n < this->maxSize in Stack and this->head < this->maxSize && this->tail < this->maxSize in Queue, as these invariants have a noticeable impact on identifying and eliminating mutants as shown in Figure 16.
This result suggests a key weakness of Daikon: it cannot differentiate between invariants that are universally true and those that hold only in the context of specific tests. We also postulate that Daikon is less mature for C++ codebases compared to Java, explaining the lack of high-quality class invariants on our benchmark. Moreover, our findings indicate that LLMs are better at capturing the concept of true ”global” or ”class” invariants, reflecting properties that are inherent to the data structure rather than incidental to the tests used.
5. Case study: Z3 bdd_manager class
Table 7. Statistics of the studied data structures in Z3
ema dlist heap hashtable permutation scoped_vector bdd_manager
# LoC 57 243 309 761 177 220 1635
# dependencies 0 2 1 9 2 2 13
As a real-world case study, we apply ClassInvGen to synthesize invariants for 7 core data structures from Z3 (de Moura and Bjørner, 2008), ranging from the simple 57-line ema class to the complex 1635-line bdd_manager. The complete set includes dlist, heap, hashtable, permutation, and scoped_vector, with varying implementation complexity and the number of dependent classes as shown in Table 7. Our results were validated by one of the Z3 authors, who confirmed at least one correct and useful invariant for each studied class, with the bdd_manager class yielding 11 valuable invariants including the 2 already present from 8 generations.
Z3 is a widely adopted SMT solver used in a variety of high-stakes applications requiring rigorous correctness, such as formal verification, program analysis, and automated reasoning. It is integrated into tools like LLVM (Lattner and Adve, 2004), KLEE (Cadar et al., 2008), Dafny (Leino, 2010) and Frama-C (Kirchner et al., 2015). We selected the Z3 codebase due to its stringent correctness requirements; as an SMT solver, Z3 is employed in applications demanding high reliability. This high-stakes environment makes Z3 an ideal testbed for assessing the effectiveness of synthesized invariants.
The bdd_manager class [^2] https://github.com/Z3Prover/z3/blob/master/src/math/dd/dd_bdd.h is particularly noteworthy. It was chosen because it is a self-contained example with developer-written unit tests for validation, presenting a realistic yet manageable challenge. Note that the existing developer tests were used after invariants were generated, not as input to the LLM. The bdd_manager class in Z3 is a utility for managing Binary Decision Diagrams (BDDs), which are data structures used to represent Boolean functions efficiently. In BDDs, Boolean functions are represented as directed acyclic graphs, where each non-terminal node corresponds to a Boolean variable, and edges represent the truth values of these variables (true or false). This representation simplifies complex Boolean expressions and enables efficient operations on Boolean functions.
With 382 lines of code in its header and 1253 lines in the implementation file, bdd_manager surpasses standard data structure complexity, offering an opportunity to evaluate ClassInvGen’s capability to generate meaningful invariants relevant to real-world scenarios. ClassInvGen achieves this by compositional generation, recursively traversing the source program’s AST (Section 3.1.2). Recursive generation became crucial when handling large classes like bdd_manager, which exceeded the LLM’s context window. Decomposing and processing its components separately allowed us to fit relevant parts into the model’s input, demonstrating the utility of recursive invariant generation for large codebases. This supports its relevance in real-world applications beyond the benchmarks.
The bdd_manager class includes a developer-written member function for checking its well-formedness, as shown in Figure 17, which we removed during ClassInvGen generation. Of the 56 invariants generated by ClassInvGen, one of Z3 main authors identified 11 distinct correct and useful invariants (e.g., Figure 18) including the 2 developer-written invariants; these invariants could potentially be integrated into the codebase. An additional 5 distinct ok invariants (e.g., Figure 19) are labeled correct but have limited utility, 16 distinct correct but useless invariants (e.g., those already checked during compilation, such as type checks and constants, Figure 20), and 2 incorrect invariants (e.g., Figure 21). The remaining invariants were repetitions within these categories. This evaluation aligns with ClassInvGen’s validation results, as our validation pipeline also identified 2 incorrect invariants that failed bdd_manager unit tests.
Overall, the Z3 authors’ evaluation results further confirm ClassInvGen’s potential utility in real-world, large-scale codebases.
1bool bdd_manager::well_formed() {
2 bool ok = true;
3 for (unsigned n : m_free_nodes) {
4 ok &= (lo(n) == 0 && hi(n) == 0 && m_nodes[n].m_refcount == 0);
5 if (!ok) {
[…]
7 display(verbose_stream()););
8 UNREACHABLE();
9 return false;
10 }
11 }
12
13 for (bdd_node const& n : m_nodes) {
14 if (n.is_internal()) continue;
15 unsigned lvl = n.m_level;
16 BDD lo = n.m_lo;
17 BDD hi = n.m_hi;
18 ok &= is_const(lo) $|$ level(lo) < lvl;
19 ok &= is_const(hi) $|$ level(hi) < lvl;
20 ok &= is_const(lo) $|$ !m_nodes[lo].is_internal();
21 ok &= is_const(hi) $|$ !m_nodes[hi].is_internal();
22 if (!ok) {
[…]
24 UNREACHABLE();
25 return false;
26 }
27 }
28 return ok;
29}
Figure 17. Z3 developer-written class invariants for bdd_manager class
1// Node consistency: Each node’s index should match its position in m_nodes
2for (unsigned i = 0; i < m_nodes.size(); ++i) {
3 assert(m_nodes[i].m_index == i);
4}
Figure 18. Correct and useful invariant for bdd_manager class
1// Cache consistency: Entries in the operation cache should be valid
2for (const auto* e : m_op_cache) {
3 assert(e != nullptr);
4 assert(e->m_result != null_bdd);
5}
Figure 19. Ok invariant for bdd_manager class
1// m_is_new_node is a boolean
2assert(m_is_new_node == true $|$ m_is_new_node == false);
Figure 20. Correct and useless invariant for bdd_manager class
1// The number of nodes should not exceed the maximum number of BDD nodes
2assert(m_nodes.size() <= m_max_num_bdd_nodes);
Figure 21. Incorrect invariant for bdd_manager class
6. Related Work
In this section, we discuss how ClassInvGen relates to previous works on synthesizing program invariants statically, dynamically and neurally.
6.1. Static approaches
Static techniques, such as interpolation (McMillan, 2004) or abstract interpretation (Cousot and Cousot, 1977) perform a symbolic analysis of source code to compute static over-approximations of runtime behavior and represent them as program invariants over suitable domains. These techniques are often used to prove the safety properties of the code. They focus on synthesizing loop invariants and method pre/postconditions, and a few around module-level specifications (Lahiri et al., 2009). Given the undecidability of program verification, these techniques scale poorly for real-world programs, especially in the presence of complex data structures and frameworks. In contrast, ClassInvGen can be applied to large codebases to synthesize high-quality class invariants but does not guarantee soundness by construction.
6.2. Dynamic approaches
Dynamic synthesis techniques, such as Daikon (Ernst et al., 2007), DIG (Nguyen et al., 2012), SLING (Le et al., 2019), and specification mining (Ammons et al., 2002), learn invariants by observing the dynamic behaviors of programs over a set of concrete execution traces. One advantage of these dynamic techniques is that they can be agnostic to the code and generally applicable to different languages. However, these approaches are limited by the templates or patterns over which the invariants can be expressed. DySy (Csallner et al., 2008) employs dynamic symbolic execution to alleviate the problem of fixed templates for bounded executions but resorts to ad-hoc abstraction for loops or recursion. Hellendoorn et al. (2019) trained models to predict the quality of invariants generated by tools such as Daikon, but do not generate new invariants. SpecFuzzer (Molina et al., 2022) generates numerous candidate assertions via fuzzing to construct templates and filters them using Daikon and mutation testing. Finally, Geminus (Boockmann and Lüttgen, 2024) aims at synthesizing sound and complete class invariants representing the set of reachable states, guiding their search using random test cases termed Random Walk.
Unlike these approaches, ClassInvGen can generate a much larger class of invariants, leveraging multimodal inputs, including source code, test cases, comments, and even the naming convention learned from training data, to enhance invariant synthesis. Further, unlike prior dynamic approaches, LLM-based test generation (an active area of research (Lemieux et al., 2023; Schäfer et al., 2023; Yang et al., 2024)) reduces the need to have a high-quality test suite to obtain the invariants.
For the use case of static verification, learning-based approaches have been used to iteratively improve the quality of the synthesized inductive invariants (Garg et al., 2014, 2016; Padhi et al., 2016) from dynamic traces. However, these approaches have not been evaluated in real-world programs due to the need for symbolic reasoning.
6.3. Neural approaches
LLM-based invariant synthesis is an emerging area of research with some noteworthy recent contributions. Pei et al. (2023) trained a model for zero-shot invariant synthesis, which incurs high training costs and lacks feedback-driven repair. Their approach uses Daikon-generated invariants as both training data and ground truth, which can lead to spurious invariants.
Prior work on nl2postcond (Endres et al., 2024) prompts LLMs to generate pre and postcondition of Python and Java benchmarks, illustrating LLMs’ ability to generate high-quality specifications. However, they do not prune incorrect invariants and do not generate class invariants that ClassInvGen does. It is an interesting future work to combine this work with ClassInvGen to generate complete class-level specifications including pre and postconditions for the public methods of the class.
For static verification, recent works include the use of LLM for intent-formalization from natural language (Lahiri, 2024), and inferring specifications and inductive program invariants (Kamath et al., 2023; Ma et al., 2024). None of these techniques scale to real-world programs due to the need for complex symbolic reasoning.
7. Limitation
ClassInvGen uses generated tests for invariant pruning, but the test suite may include spurious tests that can incorrectly prune valid invariants. The generated tests might not represent valid sequences of method calls; for example, invoking a pop() method before a push() method could fail certain assertions, leading to improper pruning.
Another limitation is the LLM’s context window, which restricts the amount of code that can be processed in a single call. This limitation makes it challenging to handle large codebases. ClassInvGen partially addresses this issue through compositional generation, breaking down the code into manageable parts. Ongoing advancements in LLMs, as highlighted in recent work (Liu et al., 2024; Gao et al., 2023), are also expected to mitigate this limitation.
For future work, we plan to integrate invariant generation with the generation of formal specifications for member functions, enabling LLM a more comprehensive understanding of program behavior. Additionally, we aim to evaluate ClassInvGen on larger and more complex systems beyond Z3, demonstrating its scalability to diverse codebases.
8. Conclusion
In this paper, we describe an approach to leverage LLMs and a lightweight mixed static/dynamic approach to synthesize class invariants. Our experiments on standard C++ data structures as well as a popular and high-assurance codebase demonstrate the feasibility of our approach. Our technique is currently limited by an automated way to integrate the generated tests into the build system of the underlying repo, and the need for developers to validate the invariants. We envision that integrating ClassInvGen with the continuous integration (CI) and pull requests (PR) can aid in scaling the approach to more developers. In future work, we also plan to investigate incorporating developer feedback to repair or strengthen generated invariants.
Acknowledgments
We would like to thank Nikolaj Bjørner for his encouragement and help in inspecting the several generated invariants for the Z3 source code and provide feedback about their utility.
Appendix A Prompt
1You are an expert in creating program invariants from code and natural language.
2Invariants are assertions on the variables in scope that hold true at different program points
3We are interested in finding invariants that hold at both start and end of a function within a data structure. Such an invariant is commonly known as an object invariant.
4
5The invariants can usually be expressed as a check on the state at the particular program point. The check should be expressed as a check in the same underlying programming language which evaluates to true or false. To express these, you can use:
6- An assertion in the programming language
7- A pure method (which does not have any side effect on the variables in scope) that checks one or more assertion
8- For a collection, you can use a loop to iterate over elements of the collection and assert something on each element or a pair of elements.
9
10Task Description:
11Task 1: Given a module, in the form of a class definition, your task is to infer object invariants about the class. For doing so, you may examine how the methods of the class read and modify the various fields of the class.
12For coming up with invariants, you may use the provided code and any comments in the code. You may also use world knowledge to guide the search for invariants.
13
14Task 2: Generate unit tests for the class based on the class definition and public API methods. The test cases should simulate a series of public method calls to verify the behavior of the class, but do not use any testing framework like gtest. Do not add assert or any form of assertions.
Figure 22. ClassInvGen Generation system prompt: instruction and task description.
15Input Format:
16You will be given the name of a class or typedef, and a section of code containing the definition of the class. You will also be given the definitions of functions that read and modify the fields of the class.
17
18Output Format:
19The output should be in the following format:
20
21The first paragraph should begin with "REASONING:". From the next line onwards, it should contain the detailed reasoning and analysis used for the inference of the object invariants. The entire text should be enclosed in \$\$\$. For example,
22‘‘\$\$\$
23REASONING:
24explanation
25\$\$\$
26
27The next paragraph should begin with "INVARIANTS:". From the next line onwards, it should contain a list of the various invariants inferred. The invariants should be in the form of code in the same underlying programming language, enclosed by ‘‘‘. Each invariant should start from a new line, and be separated by "---". Use lambda if necessary. If lambda is recursive, explicitly specify the type of the lambda function and use std::function for recursion. Do not use helper functions. For example,
28INVARIANTS:
29‘‘‘/* Invariant 1 */‘‘‘
30---
31‘‘‘/* Multi line Invariant 2 */
32 assert(...); ‘‘‘
33---
34‘‘‘/* Invariant 3 */‘‘‘
35
36The next paragraph should begin with "TESTS:". From the next line onwards, it should contain a list of a API call sequence in the form of code enclosed by ‘‘‘. Each test should start from a new line, and be separated by "---". For example,
37TESTS:
38‘‘‘/* Test 1 */
39 this->method1();
40 this->method2();‘‘‘
41---
42‘‘‘/* Test 2 */
43 this.method3(...); ‘‘‘
44---
45‘‘‘/* Test 3 */‘‘‘
46
47Important:
481. Follow the output format strictly, particularly enclosing each invariant in triple-ticks (‘‘‘), and enclosing the reasoning in \$\$\$.
492. Only find object invariants for the target class provided to you, do not infer invariants for any other class.
503. Make sure the invariant is a statement in the same underlying programming language as the source program.
514. If you can decompose a single invariant into smaller ones, try to output multiple invariants.
Figure 23. ClassInvGen Generation system prompt: input-output format.
Figure 22 and Figure 23 show the system prompt used by ClassInvGen for invariant-test co-generation. The former presents the instruction and task description, while the latter illustrates the input-output format.
1Name of Data Structure to Annotate: {struct}
2Code:
3‘‘‘
4{code}
5‘‘‘
Figure 24. ClassInvGen Generation user prompt template.
1You are an expert in repairing program invariants from code and natural language.
2Invariants are assertions on the variables in scope that hold true at different program points.
3
4We are interested in finding invariants that hold at both start and end of a function within a data structure. Such an invariant is commonly known as an object invariant.
5
6The invariants can usually be expressed as a check on the state at the particular program point. The check should be expressed as a check in the same underlying programming language which evaluates to true or false. To express these, you can use:
7- An assertion in the programming language
8- A pure method (which does not have any side effect on the variables in scope) that checks one or more assertion
9- For a collection, you can use a loop to iterate over elements of the collection and assert something on each element or a pair of elements.
10
11Task Description:
12Given a module, in the form of a class definition, your task is to infer object invariants about the class. For doing so, you may examine how the methods of the class read and modify the various fields of the class.
13
14For coming up with invariants, you may use the provided code and any comments in the code. You may also use world knowledge to guide the search for invariants.
15
16Input Format:
17You will be given the name of a class or typedef, and a section of code containing the definition of the class. You will also be given the definitions of functions which read and modify the fields of the class.
18
19Output Format:
20The output should be in the following format:
21
22The first paragraph should begin with "REASONING:". From the next line onwards, it should contain the detailed reasoning and analysis used for the inference of the object invariants. The entire text should be enclosed in \$\$\$. For example,
23‘‘\$\$\$
24REASONING:
25explanation
26\$\$\$‘‘
27
28The next paragraph should begin with "INVARIANTS:". From the next line onwards, it should contain a list of the various invariants inferred. The invariants should be in the form of code in the same underlying programming language, enclosed by ‘‘‘. Each invariant should start from a new line, and be separated by "---". For example,
29INVARIANTS:
30‘‘‘/* Invariant 1 */‘‘‘
31---
32‘‘‘/* Multi line Invariant 2 */
33 assert(...);‘‘‘
34---
35‘‘‘/* Invariant 3 */‘‘‘
36
37Important:
381. Follow the output format strictly, particularly enclosing each invariant in triple-ticks (‘‘‘), and enclosing the reasoning in \$\$\$.
392. Only find object invariants for the target class provided to you, do not infer invariants for any other class.
403. Make sure the invariant is a statement in the same underlying programming language as the source program.
414. If you can decompose a single invariant into smaller ones, try to output multiple invariants.
Figure 25. ClassInvGen Refinement system prompt: instruction and task description.
1Please fix the failed invariants given the feedback, tests and the original source code.
2
3Failed Invariant:
4‘‘‘
5{invariant}
6‘‘‘
7
8Error message for the failed invariant:
9‘‘‘
10{feedback}
11‘‘‘
12
13Name of Data Structure to Annotate: {struct}
14
15Original Code:
16‘‘‘
17{code}
18‘‘‘
19
20Gold Tests that Fail the Invariant:
21{tests}
Figure 26. ClassInvGen Refinement user prompt template.

### v2

@@
 ClassInvGen: Class Invariant Synthesis using Large Language Models
-Abstract.
+Abstract
 Formal program specifications in the form of preconditions, postconditions, and class invariants have several benefits for the construction and maintenance of programs.
@@
 In this work, we describe ClassInvGen, a method for co-generating executable class invariants and test inputs to produce high-quality class invariants for a mainstream language such as C++, leveraging LLMs’ ability to synthesize pure functions.
-We show that ClassInvGen outperforms a pure LLM-based technique to generate specifications (from code) as well as prior data-driven invariant inference techniques such as Daikon.
+We demonstrate that ClassInvGen outperforms a pure LLM-based technique for generating specifications (from code) as well as prior data-driven invariant inference techniques such as Daikon.
 We contribute a benchmark of standard C++ data structures along with a harness that can help measure both the correctness and completeness of generated specifications using tests and mutants.
 We also demonstrate its applicability to real-world code by performing a case study on several classes within a widely used and high-integrity C++ codebase.
-journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
-1. Introduction
+Keywords:
+Program Synthesis Large Language Models Class Invariants Formal Verification
+1 Introduction
 Invariants are predicates that hold on the program state for all executions of the program.
@@
 Further, for stateful classes, class invariants are facts that hold as both preconditions and postconditions of the public methods of the class, in addition to serving as a postcondition for the class constructors for the class.
-These program invariants help make explicit the assumptions on the rest of the code helping modular review, reasoning, and analysis.
+These program invariants help make explicit the assumptions on the rest of the code, helping modular review, reasoning, and analysis.
 Program invariants are useful for several aspects of software construction and maintenance during the lifetime of a program.
-First, executable program invariants can be enforced at runtime, thus serving as an early indicator of state corruption helping root causing and stopping a program instead of producing unexpected values.
-Runtime invariants serve as additional test oracles to amplify testing efforts to catch subtle bugs related to state corruption; this in turn helps with regression testing as the program evolves to satisfy new requirements.
-The utility of program invariants has led to design-by-contract in languages such as Eiffel (Meyer, 1992), as well as support in other languages such as Java (JML (Burdy et al., 2005)) and .NET (Code Contracts (Fähndrich, 2010)).
-Further, for languages that support static formal verification (e.g., Dafny (Leino, 2010), Verus (Lattuada et al., 2023), F* (Swamy et al., 2011), Frama-C (Kirchner et al., 2015)), these invariants can serve as the specification as well as help make formal verification scale by making verification modular and scalable.
+First, executable program invariants can be enforced at runtime, where they provide an early indicator of state corruption, help with root causing, and allow a program to halt with an error instead of producing unexpected values.
+Runtime invariants serve as additional test oracles to amplify testing efforts to catch subtle bugs related to state corruption; this, in turn, helps with regression testing as the program evolves to satisfy new requirements.
+The utility of program invariants has led to design-by-contract in languages such as Eiffel [32], as well as support in other languages such as Java (JML [5]) and .NET (Code Contracts [12]).
+Furthermore, for languages that support static formal verification (e.g., Dafny [26], Verus [24], F* [41], Frama-C [20]), invariants can serve as a part of the specification, helping make formal verification modular and scalable.
 Unfortunately, invariants are underutilized because they require additional work and are sometimes difficult to write, so it would be useful to find a way to generate them automatically.
 We focus specifically on automating the creation of class invariants for mainstream languages without first-class specification language support (e.g., C++) for several reasons:
-  • Class invariants are crucial to maintain the integrity of data structures and help point to state corruption that may manifest much later within the class or in the clients.
+  • Class invariants are crucial for maintaining the integrity of data structures and help point to state corruption that may manifest much later within the class or in the clients.
 Documenting such implicit contracts can greatly aid the understanding for maintainers of the class.
   • Class invariants often form important parts of preconditions and postconditions for high-integrity data structures.
-Capturing such invariants in a method and asserting them in preconditions and postconditions help reduce bloat in the specifications.
-  • Class invariants are challenging for users to write as writing them requires global reasoning across all the public methods for the class.
-For example, consider the class for a doubly linked list as implemented in Z3 Satisfiability Modulo Theories (SMT) code [^1] https://github.com/Z3Prover/z3/blob/master/src/util/dlist.h.
+Encapsulating such invariants and asserting them in preconditions and postconditions helps reduce bloat in the specifications.
+  • Class invariants are challenging for users to write, as writing them requires global reasoning across all the public methods for the class.
+For example, consider the class in Figure 1 for a doubly-linked list, as implemented in the Z3 SMT solver. [^1] https://github.com/Z3Prover/z3/blob/master/src/util/dlist.h
 1 void insert_before(T* other) {
@@
 26 }
-Figure 1. An invariant in the doubly linked list class in Z3.
+Figure 1: An invariant in the doubly linked list class in Z3.
 We see that the invariant is repeated four times: as a precondition and postcondition for the object instance this and for other.
@@
 Synthesizing program invariants has been an active line of research, with both static and dynamic analysis-based approaches.
-Static analysis approaches based on variants of abstract interpretation (Cousot and Cousot, 1977) and interpolation (Henzinger et al., 2004) create invariants that are sound by construction.
+Static analysis approaches based on variants of abstract interpretation [7] and interpolation [18] create invariants that are sound by construction.
 However, such techniques do not readily apply to mainstream programming languages with complex language constructs or require highly specialized methods that do not scale to large modules, since the invariants need to be additionally provably inductive to be retained.
-On the other hand, Daikon (Ernst et al., 2007) and successors learn invariants dynamically by instantiating a set of templates and retaining the predicates that hold on concrete test cases.
-While applicable to any mainstream language, it is well known that Daikon generated invariants overfit the test cases and are not sound for all test cases (Polikarpova et al., 2009).
-Recent works have studied fine-tuning large language models (LLMs) to learn program invariants (Pei et al., 2023) but these methods inherit the same limitation from Daikon because their training data consists of Daikon generated invariants.
+On the other hand, Daikon [11] and successors learn invariants dynamically by instantiating a set of templates and retaining the predicates that hold on concrete test cases.
+While applicable to any mainstream language, it is well known that Daikon-generated invariants overfit the test cases and are not sound for all test cases [39].
+Recent works have studied fine-tuning large language models (LLMs), to learn program invariants [38] but these methods inherit the limitations of Daikon because their training data consists of Daikon-generated invariants.
 More importantly, the approach has not been evaluated on stateful classes to construct class invariants.
-Recent work on prompting LLMs such as GPT-4 to generate program invariants for mainstream languages (Endres et al., 2024; Greiner et al., 2024; Ma et al., 2024) has been used to generate preconditions, postconditions, and loop invariants, but these methods do not readily extend to generating class invariants.
-Further, these methods cannot construct expressive invariants that require iterating over complex linked data structures (such as in Figure 1) other than simple arrays.
+Recent work on prompting LLMs such as GPT-4 to generate program invariants for mainstream languages [10, 16, 30] has been used to generate preconditions, postconditions, and loop invariants, but these methods do not readily extend to generating class invariants.
+These pipelines work at single-loop or single-method scope and validate only scalar, intraprocedural predicates, so they lack the heap-aware, cross-method perspective required to state class invariants.
+Further, these methods cannot construct expressive invariants that require iterating over complex data structures (such as in Figure 1) other than simple arrays.
 In this work, we introduce ClassInvGen, a novel method for generating high-quality object invariants for C++ classes through co-generation of invariants and test inputs using LLMs such as GPT-4o.
@@
 We leverage the framework proposed by Endres et al.
-(Endres et al., 2024) to evaluate the test-set correctness and completeness given a set of hidden validation tests and mutants.
+[10] to evaluate the test-set correctness and completeness given a set of hidden validation tests and mutants.
 We contribute a new benchmark comprising standard C++ data structures along with a harness that can help measure both the correctness and completeness of generated invariants (Section 4.1).
 We demonstrate that ClassInvGen outperforms a pure LLM-based technique for generating program invariants from code (Sections 4.34.4) as well as prior data-driven invariant inference techniques such as Daikon (Section 4.5).
-We also demonstrate its applicability for real-world code by performing a case study on ema, dlist, heap, hashtable, permutation, scoped_vector, and the most complex class, bdd_manager within the Z3 SMT solver codebase; the developers of the codebase confirmed most of the new invariants proposed by ClassInvGen for these modules (Section 5).
+We also demonstrate its applicability for real-world code by performing a case study on a set of classes in the Z3 SMT solver codebase, including the relatively complex bdd_manager class; the developers of the codebase confirmed most of the new invariants proposed by ClassInvGen for these modules (Section 5).
 Our contributions are summarized below:
-  • Present new techniques of invariant-test co-generation by combining simple static analysis with LLM and implement an end-to-end prototype (Section 3).
-  • Introduce a high-quality ClassInvGen-instrumented benchmark for evaluating object invariants (Section 4.1).
-  • Investigate LLM-assisted class invariant synthesis (Sections 4.34.5).
-  • Conduct a case study on Z3 bdd_manager module using ClassInvGen (Section 5).
-2. Running Example: AVL Tree
-Throughout this paper, AvlTree (Wikipedia contributors, 2024; Dimitrios, 2023) from our benchmark (Section 4.1) will be used as a running example to illustrate the workflow.
+  • We introduce a new technique for invariant-test co-generation by combining simple static analysis with LLMs and implement an end-to-end prototype (Section 3).
+  • We introduce a high-quality ClassInvGen-instrumented benchmark for evaluating object invariants (Section 4.1).
+  • We investigate LLM-assisted class invariant synthesis (Sections 4.34.5).
+  • We conduct a case study on Z3 class modules using ClassInvGen (Section 5).
+ClassInvGen is conceived as a specification-drafting aid: it produces candidate invariants that a developer can accept, refine, or discard, thereby following the long-advocated “human-in-the-loop” paradigm in specification mining [35].
+Our aim is not to replace expert judgement, but to accelerate it.
+2 Running Example: AVL Tree
+Throughout this paper, AvlTree [43, 9] from our benchmark (Section 4.1) will be used as a running example to illustrate the workflow.
 An AVL tree is a self-balancing binary search tree (BST) where the difference in heights between the left and right subtrees of any node (called the balance factor) is at most 1.
 This ensures that the tree remains approximately balanced.
-In this implementation shown in Figure 2, the AvlTree class contains several public methods: insert, remove, contains, clear, height, size, empty, in_order_traversal, pre_order_traversal, post_order_traversal.
+In this implementation shown in Figure 3 in Section 3.1, the AvlTree class contains several public methods: insert, remove, contains, clear, height, size, empty, in_order_traversal, pre_order_traversal, post_order_traversal.
 AvlTree maintains several class invariants determined by the authors of this paper:
@@
 The task is to infer these high-quality invariants from the source code.
+3 Approach
+Figure 2: Overview of ClassInvGen.
+An overview of the ClassInvGen framework is shown in Figure 2.
+It outlines an automated pipeline for inferring class invariants from source code.
+ClassInvGen takes a complete source program as input and outputs invariant candidates it has identified with high confidence (called filtered invariants).
+ClassInvGen starts with a preprocessing step which performs static analysis on the program (Section 3.1).
+Next, an LLM is used to generate candidate invariants and filtering test suites (Section 3.1).
+Then, the code is instrumented to facilitate checking candidate invariants (Section 3.1).
+Finally, ClassInvGen uses generated tests to prune invariant candidates (Section 3.2), and a refinement loop is used to iteratively improve the results (Section 3.3).
+3.1 Generation
+3.1.1 Preprocessing
+As illustrated in Figure 2, the generation phase begins with a static analysis of the source program.
+ClassInvGen uses a Tree-Sitter-based parser for program preprocessing; Tree-Sitter [3] is a parser generator tool that constructs a syntax tree from source files.
+ClassInvGen parses the entire source program into an abstract syntax tree (AST) to extract class members and their recursive dependencies.
+It then identifies the target class and gathers details (e.g., method declarations, field declarations, and subclass definitions) relevant to forming correct class invariants.
+ClassInvGen recursively analyzes all identified classes (i.e., the target class and its subclasses) and performs a topological sort to prepare generation from the leaf class upward as shown in Algorithm 1.
+1
+2
+3
+4Function GenerateInvariant(target_class):
+5       if target_class.id $\in$ invariants_dict then
+6             return invariants_dict[target_class.id]
+7
+8
+9      dep_classes $\leftarrow$ getClassRecursively(target_class)
+10       rev_topsorted_classes $\leftarrow$ reverseToplogicalSort(dep_classes)
+11
+12      foreach class $\in$ rev_topsorted_classes do
+13             class_code $\leftarrow$ getCodeForClass(class, invariants_dict)
+14
+15            foreach dep $\in$ getClassDependencies(class) do
+16                   // Invariant: invariants have been generated for dep
+17                   dep_code $\leftarrow$ getCodeForClass(dep, invariants_dict)
+18                   class_code $\leftarrow$ class_code + dep_code
+19
+20
+21            invariants_dict[class.id] $\leftarrow$ generateInvariantWithLLM(class_code)
+22
+23
+24      return invariants_dict[target_class.id]
+25
+26
+27Function getCodeForClass(class, invariants_dict):
+28       class_text $\leftarrow$ class.get_declaration_text()
+29        // Header
+30       if class.id $\in$ invariants_dict and invariants_dict[class.id] then
+31             // Get generated invariants
+32             class_text $\leftarrow$ invariants_dict[class.id] + class_text
+33
+34      return class_text
+35
+Algorithm 1 Function to generate invariants for target class AST
+3.1.2 Generation by LLM
+After building the source program AST, ClassInvGen uses LLMs to analyze the class module and infers both invariants for the target class and tests that exercise the class’s implementation as thoroughly as possible.
+ClassInvGen uses a fixed system prompt that defines class invariants and outlines two main tasks: (1) generating class invariants from the source code, and (2) creating a test suite of valid API calls without specifying expected outputs (see prompt details in the Appendix 0.A).
+Next, ClassInvGen instantiates a user prompt template with the actual target class.
+From the source program AST, ClassInvGen identifies program dependencies and populates the prompt template with the leaf struct/class.
+Starting from the leaf nodes, ClassInvGen leverages previously generated invariants by including them in the prompt for later classes.
+To accommodate the LLM’s context window limit, only the relevant child classes of the current target class are included in the prompt, with method implementations and private fields/methods hidden when necessary.
+An algorithm for this process is presented in Algorithm 1.
+Algorithm 1 presents the invariant generation process for a source program AST.
+The main function GenerateInvariant takes a target_class and leverages a caching mechanism through invariants_dict to avoid redundant computations (Line 1).The algorithm first collects dependent classes via getClassRecursively (Line 1) and sorts them using reverseToplogicalSort to ensure dependency-aware processing (Line 1).
+For each class in the sorted order, it constructs the necessary context by obtaining the class code through getCodeForClass (Line 1).
+For each dependency dep of the current class, it retrieves the dep_code and concatenates it with class_code (Lines 11).
+The algorithm then generates invariants using generateInvariantWithLLM and stores them in invariants_dict (Line 1).
+The helper function getCodeForClass constructs class representations by combining the declaration text with any existing invariants from invariants_dict (Lines 11).
+The algorithm concludes by returning final_invariants for the target class, effectively managing the invariant generation process while respecting LLM context limitations.
+ClassInvGen accommodates large codebases by dividing the source program into smaller modules that fit within the LLM’s context window.
+It then iteratively generates invariants and test cases, starting from leaf classes and working up towards the root class.
+At each step, ClassInvGen leverages previous invariants generated for child classes to inform the invariants for parent classes.
+For the AvlTree example, we begin by instantiating the prompt with Node for annotation, followed by AvlTree, since Node is a subclass of AvlTree, as illustrated in Figure 3.
+In this specific case, however, Algorithm 1 does not make a difference due to the small size of the source program; the entire AvlTree code fits within the LLM’s context window easily.
 1class AvlTree {
@@
 26}
-Figure 2. Header file of AvlTree.
-3. Approach
-Figure 3. Overview of ClassInvGen.
-An overview of the ClassInvGen framework is shown in Figure 3.
-It outlines an automated pipeline for inferring class invariants from source code.
-ClassInvGen takes a complete source program as input and outputs invariant candidates it has identified with high confidence (called filtered invariants).
-ClassInvGen starts with a preprocessing step which performs static analysis on the program (Section 3.1.1).
-Next, an LLM is used to generate candidate invariants and filtering test suites (Section 3.1.2).
-Then, the code is instrumented to facilitate checking candidate invariants (Section 3.1.3).
-Finally, ClassInvGen uses generated tests to prune invariant candidates (Section 3.2), and a refinement loop is used to iteratively improve the results (Section 3.3).
-3.1. Generation
-3.1.1. Preprocessing
-As illustrated in Figure 3, the generation phase begins with a static analysis of the source program.
-ClassInvGen uses a Tree-Sitter-based parser for program preprocessing; Tree-Sitter (Brunsfeld et al., 2024) is a parser generator tool that constructs a syntax tree from source files.
-ClassInvGen parses the entire source program into an abstract syntax tree (AST) to extract class members and their recursive dependencies.
-It then identifies the target class and gathers details (e.g., method declarations, field declarations, and subclass definitions) relevant to forming correct class invariants.
-ClassInvGen recursively analyzes all identified classes (i.e., the target class and its subclasses) and performs a topological sort to prepare generation from the leaf class upward as shown in Algorithm 1.
-1
-2
-3
-4Function GenerateInvariant(target_class):
-5       if target_class.id $\in$ invariants_dict then
-6             return invariants_dict[target_class.id]
-7
-8
-9      dep_classes $\leftarrow$ getClassRecursively(target_class)
-10       rev_topsorted_classes $\leftarrow$ reverseToplogicalSort(dep_classes)
-11
-12      foreach class $\in$ rev_topsorted_classes do
-13
-14            class_code $\leftarrow$ getCodeForClass(class, invariants_dict)
-15
-16            foreach dep $\in$ getClassDependencies(class) do
-17                   // Invariant: dep has been generated invariants for
-18                   dep_code $\leftarrow$ getCodeForClass(dep, invariants_dict)
-19                   class_code $\leftarrow$ class_code + dep_code
-20
-21
-22            invariants_dict[class.id] $\leftarrow$ generateInvariantWithLLM(class_code)
-23
-24
-25      final_invariants $\leftarrow$ invariants_dict[target_class.id]
-26       return final_invariants
-27
-28
-29Function getCodeForClass(class, invariants_dict, include_method_bodies=False):
-30       class_text $\leftarrow$ class.get_declaration_text()
-31        // Header code
-32       if class.id $\in$ invariants_dict and invariants_dict[class.id] then
-33             class_text $\leftarrow$ invariants_dict[class.id] + class_text
-34              // Get generated invariants
-35
-36
-37      if include_method_bodies then
-38             // Add methods bodies to class_text if context allows
-39
-40      return class_text
-41
-Algorithm 1 Function to generate invariants for target class AST parsed by Tree-Sitter
-3.1.2. Generation by LLM
-1You are an expert in creating program invariants from code and natural language.
-2Invariants are assertions on the variables in scope that hold true at different program points
-3We are interested in finding invariants that hold at both start and end of a function within a data structure. Such an invariant is commonly known as an object invariant.
-4
-5The invariants can usually be expressed as a check on the state at the particular program point.
-The check should be expressed as a check in the same underlying programming language which evaluates to true or false.
-To express these, you can use:
-6- An assertion in the programming language
-7- A pure method (which does not have any side effect on the variables in scope) that checks one or more assertion
-8- For a collection, you can use a loop to iterate over elements of the collection and assert something on each element or a pair of elements.
-9
-10Task Description:
-11Task 1: Given a module, in the form of a class definition, your task is to infer object invariants about the class.
-For doing so, you may examine how the methods of the class read and modify the various fields of the class.
-12For coming up with invariants, you may use the provided code and any comments in the code. You may also use world knowledge to guide the search for invariants.
-13
-14Task 2: Generate unit tests for the class based on the class definition and public API methods.
-The test cases should simulate a series of public method calls to verify the behavior of the class, but do not use any testing framework like gtest.
-Do not add assert or any form of assertions.
-Figure 4. ClassInvGen Generation system prompt: instruction and task description.
-After building the source program AST, ClassInvGen uses LLMs to analyze the class module and infers both invariants for the target class and tests that exercise the class’s implementation as thoroughly as possible.
-ClassInvGen uses a fixed system prompt that defines class invariants and outlines two main tasks: (1) generating class invariants from the source code, and (2) creating a test suite of valid API calls without specifying expected outputs (Figure 4 and Figure 23 in Appendix).
-Next, ClassInvGen instantiates a user prompt template (Figure 24 in Appendix) with the actual target class.
-From the source program AST, ClassInvGen identifies program dependencies and populates the prompt template with the leaf struct/class.
-Starting from the leaf nodes, ClassInvGen leverages previously generated invariants by including them in the prompt for later classes.
-To accommodate the LLM’s context window limit, only the relevant child classes of the current target class are included in the prompt, with method implementations and private fields/methods hidden when necessary.
-An algorithm for this process is presented in Algorithm 1.
-Algorithm 1 presents the invariant generation process for a source program AST.
-The main function GenerateInvariant takes a target_class and leverages a caching mechanism through invariants_dict to avoid redundant computations (Line 1).
-The algorithm first collects dependent classes via getClassRecursively (Line 1) and sorts them using reverseToplogicalSort to ensure dependency-aware processing (Line 1).
-For each class in the sorted order, it constructs the necessary context by obtaining the class code through getCodeForClass (Line 1).
-For each dependency dep of the current class, it retrieves the dep_code and concatenates it with class_code (Lines 11).
-The algorithm then generates invariants using generateInvariantWithLLM and stores them in invariants_dict (Line 1).
-The helper function getCodeForClass constructs class representations by combining the declaration text with any existing invariants from invariants_dict (Lines 11).
-It optionally includes method bodies based on context window constraints.
-This approach ensures efficient invariant generation while maintaining all necessary context and dependencies (Line 1).
-The algorithm concludes by returning final_invariants for the target class, effectively managing the invariant generation process while respecting LLM context limitations.
-ClassInvGen accommodates large codebases by dividing the source program into smaller modules that fit within the LLM’s context window.
-It then iteratively generates invariants and test cases, starting from leaf classes and working up towards the root class.
-At each step, ClassInvGen leverages previous invariants generated for child classes to inform the invariants for parent classes.
-For the AvlTree example, we begin by instantiating the prompt with Node for annotation, followed by AvlTree, since Node is a subclass of AvlTree, as illustrated in Figure 2.
-In this specific case, however, Algorithm 1 does not make a difference due to the small size of the source program; the entire AvlTree code fits within the LLM’s context window easily.
+Figure 3: Header file of AvlTree.
 In contrast, when working with the bdd_manager class in Z3 (around 1700 lines of code), ClassInvGen begins generation with bdd, a subclass of bdd_manager.
-Algorithm 1 enables ClassInvGen to partition bdd_manager class and outputs meaningful class invariants (see Section 5).
-3.1.3. Instrumentation
-1bool AvlTree::empty() {
-2 check_invariant();
-3 auto ret = empty_original();
-4 check_invariant();
-5 return ret;
-6}
-7
-8bool AvlTree::empty_original() { return n == 0; }
-Figure 5. AvlTree Instrumented with invariants
-1void AvlTree::check_invariant() {
-2 std::function<bool(const std::unique_ptr<Node>&)> is_balanced = [&](const std::unique_ptr<Node>& node) -> bool {
-3 if (!node) return true;
-4 int left_height = height(node->left);
-5 int right_height = height(node->right);
-6 if (std::abs(left_height - right_height) > 1) return false;
-7 return is_balanced(node->left) && is_balanced(node->right);
-8 };
-9 assert(is_balanced(root));
-10
-11}
-Figure 6. Example of a correct AvlTree class invariant
-1void AvlTree::check_invariant() {
-2 assert(height(root) == get_height(root));
-3}
-Figure 7. Example of an incorrect invariant of AvlTree because there is no get_height method
-To check candidate invariants, each public method is automatically instrumented with a check_invariant call at both the start and end of its implementation, as shown in Figure 5.
-ClassInvGen instruments these invariants to use the compiler to rule out syntactic/type-based errors in the invariants.
-Each invariant is implemented as a method call to prevent conflicts with local variables within the two invocations to check_invariant.
-To check a specific invariant, the code for the invariant is plugged in as the implementation for the check_invariant function, and assertions are added to check that the invariant holds, as depicted in Figures 6 and 7.
-This setup ensures that during pruning, described below, whenever a public API call is made, each invariant candidate is automatically checked.
-3.2. Heuristic Pruning
-As discussed above, the LLM generates test suites for candidates.
-ClassInvGen wishes to select the suite that is most effective at identifying inaccurate invariants.
-To select the most suitable test suite, we employ line coverage as a metric.
-Line coverage is used to select the best test suite because it serves as a straightforward proxy for test suite completeness.
-The tests with the highest coverage are considered filtering tests and are used to prune the candidate invariants.
-Notably, when generating tests, ClassInvGen focuses on creating valid sequences of API calls to exercise the class under test but does not assert the expected behavior.
-This approach simplifies the process by eliminating the need for a test oracle, which is unnecessary in the ClassInvGen framework since the purpose of test generation is to filter invariant candidates, not to test the source program directly.
+Algorithm 1 enables ClassInvGen to partition the bdd_manager class and outputs meaningful class invariants (see Section 5).
+3.1.3 Instrumentation
+To check candidate invariants, each public method is automatically instrumented with a check_invariant call at both the start and end of its implementation.
+This allows us to verify that invariants hold both before and after method execution.
+Each invariant is implemented as a method call to prevent conflicts with local variables.
+When a specific invariant is being checked, its code is plugged into the check_invariant function with assertions.
+This ensures that during pruning, whenever a public API call is made, each invariant candidate is automatically verified.
+Additional examples of instrumentation and invariant checking are provided in Appendix 0.C.
+3.2 Heuristic Pruning
+The LLM generates test suites that serve as filters for invariant candidates.
+To select the most effective test suite, we use line coverage as a metric, as it provides a straightforward proxy for test suite completeness.
+The test suite with the highest coverage becomes our set of filtering tests.
+When generating tests, ClassInvGen creates valid sequences of API calls without asserting expected behavior, since our goal is to filter invariant candidates rather than test the source program directly.
+Among all generated test suites, we compile and run each one with the source program, selecting the one with the highest line coverage as the filtering tests.
 […]
-[…]
-These filtering tests, which are assumed to contain valid sequences, are then used to identify high-quality invariant candidates, referred to as filtered invariants.
-If an invariant candidate successfully compiles and runs with the filtering tests, it is designated a filtered invariant and the final output of ClassInvGen.
-Figure 6 illustrates a correct filtered invariant, stating that the height of an AvlTree is balanced.
-In contrast, Figure 7 shows an incorrect invariant that was filtered out by the filtering tests during compilation because it references a get_height method, which is not accessible to the user.
-To filter invariants, ClassInvGen compiles and runs candidate invariants with the filtering tests, retaining those that do not cause crashes and adding them to the set of filtered invariants.
-1int main() {
-2 // Test Case 1
-3 // Basic insertions and traversals
-4 {
-5 AvlTree tree;
-6 tree.insert(10);
-7 tree.insert(20);
-8 tree.insert(5);
-9 tree.in_order_traversal();
-10 tree.pre_order_traversal();
-11 tree.post_order_traversal();
-12 }
-13 // Test Case 2
-14 // Checking size, height, and empty status
-15 {
-16 AvlTree tree;
-17 tree.insert(10);
-18 tree.insert(20);
-19 tree.insert(5);
-20 tree.size();
-21 tree.height();
-22 tree.empty();
-23 }
-24 ... // the rest tests are skipped
-25}
-Figure 8. A test suite generated for AvlTree
-3.3. Refinement
-Invariants that fail at either compilation or runtime are candidates for refinement loops.
-For such failing invariants, ClassInvGen implements a feedback loop by collecting compiler output, error messages, and the results of filtering tests, then feeding this information back to the LLM using a dedicated prompt template (Figures 2526 in Appendix) to generate refined invariants.
-The system prompt for refinement is fixed, while the user prompt is instantiated with the failing invariant, compiler error messages, the target class’s source code, and the filtering tests that caused the failure.
-For instance, the GCC error output shown in Figure 9 for the AvlTree candidate invariant in Figure 10 leads to the refined candidate invariant in Figure 11.
-We set the default threshold for refinement loops to 3, balancing the cost of calling LLMs with the benefit of repairing a single invariant; empirically, this number works well, though it may require adjustment for different codebases.
-If the invariant still fails after 3 refinement attempts, ClassInvGen discards it and moves on.
-Invariants are accepted as filtered only after successfully passing all filtered tests.
-1avl_tree.cpp: In lambda function:
-2avl_tree.cpp:21:16: error: use of ’is_balanced’ before deduction of ’auto’
-Figure 9. gcc compiler error messages
-1void AvlTree::check_invariant() {
-2  auto is_bst = [&](const std::unique_ptr<Node>& node, const T& min, const T& max) -> bool {
-3 if (!node) return true;
-4 if (node->data <= min $|$ node->data >= max) return false;
-5 return is_bst(node->left, min, node->data) && is_bst(node->right, node->data, max);
-6 };
-7
-8 assert(is_bst(root, std::numeric_limits<T>::min(), std::numeric_limits<T>::max()));
-9}
-Figure 10. BST property before refinement: the word auto is causing the compiler to complain
-1void AvlTree::check_invariant() {
-2  std::function<bool(const std::unique_ptr<Node>&)> is_bst = [&](const std::unique_ptr<Node>& node, const T& min, const T& max) -> bool {
-3 if (!node) return true;
-4 if (node->data <= min $|$ node->data >= max) return false;
-5 return is_bst(node->left, min, node->data) && is_bst(node->right, node->data, max);
-6 };
-7
-8 assert(is_bst(root, std::numeric_limits<T>::min(), std::numeric_limits<T>::max()));
-9}
-Figure 11. BST property after refinement: auto is changed to explicit declarations. This change resolved type ambiguity, allowing the code to compile successfully.
-4. Results
-4.1. Benchmark
-Table 1.
-Characteristics of the benchmark data structures.
+If an invariant candidate successfully compiles and runs with the filtering tests, it is designated a filtered invariant and included in the final output of ClassInvGen.
+3.3 Refinement
+For invariants that fail during compilation or runtime, ClassInvGen implements a feedback-driven refinement process.
+For each failing invariant, the system collects compiler output, error messages, and test results, and then feeds this information back to the LLM using a dedicated prompt template.
+This feedback loop allows ClassInvGen to repair failing invariants by providing the LLM with specific error information and the context in which the error occurred.
+We set a default threshold of 3 refinement attempts per invariant, balancing the cost of LLM calls with the benefit of repairs.
+Refinement allows ClassInvGen to fix common issues such as type errors, undefined references, and logical inconsistencies.
+More detailed examples of the refinement process are provided in Appendix 0.C.
+4 Results
+4.1 Benchmark
+Table 1: Characteristics of the benchmark data structures.
 # LoC represents lines of code, # methods indicates the number of implemented methods, and # dep.
@@
 # dep. 1 1 0 0 0 0 1 0 0
-We use a C++ implementation of classical data structures for our micro-benchmarks (Dimitrios, 2023), which include 9 data structures with associated unit tests: AvlTree, BinarySearchTree, HashTable, Heap, LinkedList, Queue, RedBlackTree, Stack, and Vector.
+We use a C++ implementation of classical data structures for our micro-benchmarks [9], which include 9 data structures with associated unit tests: AvlTree, BinarySearchTree, HashTable, Heap, LinkedList, Queue, RedBlackTree, Stack, and Vector.
 Table 1 shows the statistics of these benchmark data structures.
@@
 All subsequent experiments are based on this benchmark setup.
-In addition to textbook examples, we also conducted experiments on utility classes (z3Prover, 2024) from Z3 (de Moura and Bjørner, 2008), including ema, dlist, heap, hashtable, permutation, scoped_vector, and the most complex class, bdd_manager.
+In addition to textbook examples, we also conducted experiments on utility classes [46] from Z3 [34], including ema, dlist, heap, hashtable, permutation, scoped_vector, and the most complex class, bdd_manager.
 The latter will be discussed in detail as a case study, including an evaluation with one of the authors of Z3.
-4.2. Evaluation
+4.2 Evaluation
 In this section, we evaluate the quality of ClassInvGen invariants. Specifically, we explore the following research questions:
-  1. (1)
-    How many of these invariants are correct (with respect to the user-provided test cases) and do they capture essential properties of the source code (Section 4.3)?
-  2. (2)
+  1. 1.
+    How many of these invariants are correct (with respect to the user-provided test cases), and do they capture essential properties of the source code (Section 4.3)?
+  2. 2.
     How complete are the invariants in their ability to distinguish the correct program from buggy counterparts (Section 4.4)?
-  3. (3)
+  3. 3.
     How does ClassInvGen compare to a state-of-the-art technique in invariant generation (namely Daikon, the most widely adopted tool for dynamic invariant synthesis) (Section 4.5)?
 Our experiments were conducted on a machine with 24 CPU cores and 64 GB of RAM. We implemented ClassInvGen using GPT-4o as the underlying LLM, with its default temperature setting of 1.
-Figure 12. Evaluation of ClassInvGen generated invariants
-4.3. Correctness
-ClassInvGen produces filtered invariants, which we evaluate using an automated pipeline (Figure 12) to assess their quality against the data structures benchmark described in Section 4.1.
-The evaluation process begins by instrumenting the source code with the candidate invariants, as detailed in Section 3.1.3.
-Unlike the generated tests used in ClassInvGen for filtering, the evaluation relies on unit tests containing input-output pairs.
-A filtered invariant is considered correct if it does not report any errors for any tests that successfully compile and run.
-Notably, these unit tests are not available during the ClassInvGen generation phase and are used exclusively for evaluation purposes.
-For each passing filtered invariant, the authors of the paper conducted a manual review.
-This examination confirmed that all validated filtered invariants are indeed correct and effectively capture properties of the data structures.
-e.g., all 3 properties of AvlTree described in Section 2 are found in ClassInvGen invariants.
+Figure 4: Evaluation of ClassInvGen generated invariants
+4.3 Correctness
+ClassInvGen produces filtered invariants, which we evaluate using an automated pipeline (Figure 4) against our benchmark (Section 4.1).
+A filtered invariant is considered correct if it reports no errors for any tests that successfully compile and run.
+Our manual review confirmed that all validated filtered invariants are indeed correct, capturing essential properties of the data structures.
 […]
@@
 […]
-# invariants # compiles # pass unit tests unit test pass rate
+Example # inv. # compiles # pass tests pass rate
 avl_tree 30 28 17 56.7%
@@
 […]
-Example # filtering tests # filtered inv. # good inv. unit tests pass rate
+Example # filter tests # filtered inv. # good inv. pass rate
 (coverage) (1 refine) (1 refine)
@@
 Average 9.44 (94.0%) 17.22 21.78 100.0%
-Table 3.
-For each example, the table shows the number and line coverage of filtering tests, the number of filtered invariants without refinement, and with 1 refinement, as well as the unit test pass rate after 1 refinement.
+Table 3: For each example, the table shows the number and line coverage of filtering tests, the number of filtered invariants without refinement, and the number of invariants after 1 refinement attempt.
 With filtering tests and 1 refinement, ClassInvGen achieves a perfect unit test pass rate.
-4.4. Completeness
-Table 4. ClassInvGen Performance Over Baseline for Previously Survived Mutants. The table shows additional mutants killed by ClassInvGen compared to the baseline and percentage improvement.
-Data Structure Unsolved Baseline (#(%)) Add. Solved by ClassInvGen (#) Improv. (%)
+4.4 Completeness
+Table 4: ClassInvGen The table shows the number of mutants that survive unit tests alone (Unsolved Base) as well as the number of additional mutants killed by ClassInvGen.
+Data Structure Unsolved Base (#(%)) Add. by ClassInvGen (#) Impr. (%)
 binary_search_tree 107(23.67) 7 6.54
@@
 Total 1057 118 11.16
-To evaluate the completeness of ClassInvGen invariants, we use a variation of mutation testing used in prior work to evaluate the strength of correct specifications (Endres et al., 2024).
-Mutation testing creates copies of the code with small changes.
-Individual copies are called mutants.
-If a mutant fails tests, the mutant is said to be killed, otherwise it survives.
-We expect that many mutations will corrupt data structures, so the number of mutants killed by ClassInvGen invariants is an indication of the strength of the invariant, or how complete those invariants are.
-Note that we only consider invariants that are correct.
-We use the freeware tool mutate_cpp (Lohmann and other contributors, 2017) to perform program mutation, producing between 236 and 682 mutants per program.
-There are three possible outcomes for each mutant: it fails to compile, it crashes during execution, or it compiles and runs successfully without errors.
-We focus exclusively on mutants that either compile successfully but crash during execution or survive execution without error, as these are distinguishable cases.
-Mutants that fail to compile remain consistent across all configurations and therefore do not contribute to meaningful differentiation.
-Figure 13. Completeness Experiment Result. The 3 bars from left to right are Tests, ClassInvGen, Tests+ClassInvGen. Tests+ClassInvGen kills the most mutants.
-We conducted experiments to evaluate the completeness of three configurations: unit tests only, ClassInvGen only, unit tests with ClassInvGen.
-The last configuration has the strongest test oracles.
-For ClassInvGen only, we removed all assertions checking the expected output in the unit tests, leaving only the sequence of method calls consistent with the original tests.
+To evaluate completeness, we use mutation testing.
+We generate mutants using mutate_cpp [29], producing between 236 and 682 mutants per program.
+We focus on mutants that either compile successfully but crash during execution or survive execution without errors.
+Figure 5: Completeness Experiment Result. The 3 bars from left to right are Tests, ClassInvGen, Tests+ClassInvGen. Tests+ClassInvGen kills the most mutants.
 […]
-Figure 13 indicates tests with ClassInvGen kill the most mutants, showing that instrumenting unit tests with ClassInvGen invariants enhances their ability to identify buggy implementations.
-Figure 14 shows HashTable a mutant that survived tests but killed by ClassInvGen invariant assert(this->_size > 0 || this->table.empty()).
-Figure 15 shows another example killed by ClassInvGen invariant assert(load_factor > 0).
+Figure 5 shows how many mutants are killed with each configuration for each benchmark.
+Combining tests with ClassInvGen kills the most mutants.
+Figures 6 and 7 show examples of mutants that survived unit tests but were killed by ClassInvGen invariants.
 1 void HashTable::clear_table() {
@@
 7}
-Figure 14. Mutant that survived unit tests but killed by ClassInvGen
+Figure 6: Mutant that survived unit tests but was killed by ClassInvGen
 1 this->hash_function = hash_function;
@@
 8 }
-Figure 15. Another Mutant that survived unit tests but killed by ClassInvGen
-However, mutation testing has its limitations.
-For example, a mutant may be equivalent to the original program (e.g., n = 0 and n &= 0), resulting in no observable behavioral difference.
-Additionally, certain mutants may fall outside the scope of the invariants to detect, such as when only observer methods are altered, which do not affect the program’s state directly.
-4.5. Comparison of ClassInvGen v.s. Daikon
-We compare ClassInvGen with Daikon, the most widely adopted dynamic invariant generator (Ernst et al., 2007).
-We replaced the invariant synthesis component in Figure 3 with Daikon, using filtering tests to generate program traces.
-We then ran Daikon’s invariant detector on these traces with a default confidence limit equal to 0.99 (larger values yield stronger filtering).
-For the most comprehensive set of invariant candidates, we combined all invariants Daikon generated at each function’s ENTER and EXIT points and used ClassInvGen invariant instrumentation to identify filtered invariants using unit tests to filter.
-The final output is a set of Daikon invariants that successfully pass unit test execution.
-On average, each benchmark example has around 5 Daikon invariants; however, in some cases, 0 to 2 of these invariants may be incorrect as shown in Table 6 examined by humans.
-It is important to note that Daikon invariants are not always correct, even with respect to the tests used to infer them.
-Daikon outputs an invariant if its confidence level exceeds a specified configuration threshold, regardless of whether the invariant holds universally.
-Through the manual review, we identified 7 incorrect class invariants, as summarized in Table 5.
-These invariants pass unit test validation because both the filtering tests and the unit tests coincidentally constructed data structures with similar parameters , producing similar program traces.
-Consequently, validation was unsuccessful because these Daikon invariants based on filtering test traces also happened to hold true on the unit tests.
-For instance, when Vector objects are constructed in both filtering tests and ground truth unit tests, elements are consistently push_backed in increasing order.
-However, users of Vector are not restricted to this behavior and could, in practice, choose to push_back larger elements first, breaking this assumed invariant.
-Therefore, this->data[] sorted by < is not a true class invariant.
-Table 5. Incorrect Daikon Invariants that pass unit tests validation.
-Data Structure Invariant Description
-HashTable this->_size one of {5, 10, 40}
-Heap this->data._Vector_base< int, std::allocator<int> >. _M_impl._Vector_impl_data._M_start[] elements >= 1
-Vector this->data[] sorted by <
-Queue this->maxSize == 10
-LinkedList this->tail[].data elements >= 1
-Stack this->maxSize one of {10, 160, 1280}
-this->maxSize == 10
-Table 6. Daikon Incorrect Invariants per Benchmark
+Figure 7: Another Mutant that survived unit tests but was killed by ClassInvGen
+4.5 Comparison of ClassInvGen vs. Daikon
+We compared ClassInvGen with Daikon [11], using filtering tests to generate program traces for Daikon’s invariant detector.
+On average, each benchmark example has around 5 Daikon invariants, with some being incorrect (see summary in Table 5 and complete results of invariants in Appendix 0.B).
+Table 5: Daikon Incorrect Invariants per Benchmark
 Data Structure Total # Invariants Incorrect Invariants
@@
 Average 5.2 0.78
-Figure 16. Daikon vs. ClassInvGen Kills
-Among the 40 correct class invariants generated by Daikon and reviewed by us, the most common invariants indicate that the class pointer is not null and remains unchanged throughout the object’s lifetime (27 invariants).
-The second most frequent type asserts that the number of elements (typically this->n) is non-negative (6 invariants).
-These two categories provide minimal utility in distinguishing and eliminating mutants (incorrect programs).
-The most valuable invariants are those that capture meaningful constraints, such as this->n < this->maxSize in Stack and this->head < this->maxSize && this->tail < this->maxSize in Queue, as these invariants have a noticeable impact on identifying and eliminating mutants as shown in Figure 16.
-This result suggests a key weakness of Daikon: it cannot differentiate between invariants that are universally true and those that hold only in the context of specific tests.
-We also postulate that Daikon is less mature for C++ codebases compared to Java, explaining the lack of high-quality class invariants on our benchmark.
-Moreover, our findings indicate that LLMs are better at capturing the concept of true ”global” or ”class” invariants, reflecting properties that are inherent to the data structure rather than incidental to the tests used.
-5. Case study: Z3 bdd_manager class
-Table 7. Statistics of the studied data structures in Z3
+Through manual review, we identified 7 incorrect Daikon invariants that pass unit test validation.
+These invariants pass because both the filtering tests and unit tests coincidentally constructed similar data structures.
+Figure 8: Daikon vs. ClassInvGen Kills
+Most Daikon invariants simply indicate that class pointers are not null (27 of 40 correct invariants) or that element counts are non-negative (6 invariants).
+The most valuable invariants, like this->n < this->maxSize in Stack, have more impact on identifying mutants (Figure 8).
+This shows a key weakness of Daikon: it cannot differentiate between universally true invariants and those that hold only in specific test contexts.
+LLMs are better at capturing true “class” invariants that are inherent to the data structure rather than incidental to the tests.
+5 Case study: Z3 bdd_manager class
+Table 6: Statistics of the studied data structures in Z3
 ema dlist heap hashtable permutation scoped_vector bdd_manager
@@
 # dependencies 0 2 1 9 2 2 13
-As a real-world case study, we apply ClassInvGen to synthesize invariants for 7 core data structures from Z3 (de Moura and Bjørner, 2008), ranging from the simple 57-line ema class to the complex 1635-line bdd_manager.
+As a real-world case study, we apply ClassInvGen to synthesize invariants for 7 core data structures from Z3 [34], ranging from the simple 57-line ema class to the complex 1635-line bdd_manager.
 The complete set includes dlist, heap, hashtable, permutation, and scoped_vector, with varying implementation complexity and the number of dependent classes as shown in Table 6.
-Our results were validated by one of the Z3 authors, who confirmed at least one correct and useful invariant for each studied class, with the bdd_manager class yielding 11 valuable invariants including the 2 already present from 8 generations.
-Z3 is a widely adopted SMT solver used in a variety of high-stakes applications requiring rigorous correctness, such as formal verification, program analysis, and automated reasoning.
-It is integrated into tools like LLVM (Lattner and Adve, 2004), KLEE (Cadar et al., 2008), Dafny (Leino, 2010) and Frama-C (Kirchner et al., 2015).
+Our results were validated by one of the Z3 authors, who confirmed at least one correct and useful invariant for each studied class, with the bdd_manager class yielding 11 valuable invariants including the 2 already written by Z3 authors.
+Z3 is a widely adopted SMT solver used in a variety of high-stakes applications that depend on its correctness.
+It is integrated into tools like LLVM [23], KLEE [6], Dafny [26] and Frama-C [20].
 We selected the Z3 codebase due to its stringent correctness requirements; as an SMT solver, Z3 is employed in applications demanding high reliability.
@@
 Note that the existing developer tests were used after invariants were generated, not as input to the LLM.
-The bdd_manager class in Z3 is a utility for managing Binary Decision Diagrams (BDDs), which are data structures used to represent Boolean functions efficiently.
+The bdd_manager class in Z3 is a utility for managing Binary Decision Diagrams (BDDs) [4], which are data structures used to represent Boolean functions efficiently.
 In BDDs, Boolean functions are represented as directed acyclic graphs, where each non-terminal node corresponds to a Boolean variable, and edges represent the truth values of these variables (true or false).
@@
 With 382 lines of code in its header and 1253 lines in the implementation file, bdd_manager surpasses standard data structure complexity, offering an opportunity to evaluate ClassInvGen’s capability to generate meaningful invariants relevant to real-world scenarios.
-ClassInvGen achieves this by compositional generation, recursively traversing the source program’s AST (Section 3.1.2).
+ClassInvGen achieves this by compositional generation, recursively traversing the source program’s AST.
 Recursive generation became crucial when handling large classes like bdd_manager, which exceeded the LLM’s context window.
@@
 This supports its relevance in real-world applications beyond the benchmarks.
-The bdd_manager class includes a developer-written member function for checking its well-formedness, as shown in Figure 17, which we removed during ClassInvGen generation.
-Of the 56 invariants generated by ClassInvGen, one of Z3 main authors identified 11 distinct correct and useful invariants (e.g., Figure 18) including the 2 developer-written invariants; these invariants could potentially be integrated into the codebase.
+The bdd_manager class has a developer-written member function for checking its well-formedness, as shown in Figure 9, which we removed during ClassInvGen generation.
+Of the 56 invariants generated by ClassInvGen, one of Z3’s main authors identified 11 distinct correct and useful invariants (e.g., Figure 10) including the 2 developer-written invariants; these invariants could potentially be integrated into the codebase.
 An additional 5 distinct ok invariants (e.g., Figure 11) are labeled correct but have limited utility, 16 distinct correct but useless invariants (e.g., those already checked during compilation, such as type checks and constants, Figure 12), and 2 incorrect invariants (e.g., Figure 13).
@@
 29}
-Figure 17. Z3 developer-written class invariants for bdd_manager class
+Figure 9: Z3 developer-written class invariants for bdd_manager class
 1// Node consistency: Each node’s index should match its position in m_nodes
@@
 4}
-Figure 18. Correct and useful invariant for bdd_manager class
+Figure 10: Correct and useful invariant for bdd_manager class
 1// Cache consistency: Entries in the operation cache should be valid
@@
 5}
-Figure 19. Ok invariant for bdd_manager class
+Figure 11: Ok invariant for bdd_manager class
 1// m_is_new_node is a boolean
 2assert(m_is_new_node == true $|$ m_is_new_node == false);
-Figure 20. Correct and useless invariant for bdd_manager class
+Figure 12: Correct and useless invariant for bdd_manager class
 1// The number of nodes should not exceed the maximum number of BDD nodes
 2assert(m_nodes.size() <= m_max_num_bdd_nodes);
-Figure 21. Incorrect invariant for bdd_manager class
-6. Related Work
-In this section, we discuss how ClassInvGen relates to previous works on synthesizing program invariants statically, dynamically and neurally.
-6.1. Static approaches
-Static techniques, such as interpolation (McMillan, 2004) or abstract interpretation (Cousot and Cousot, 1977) perform a symbolic analysis of source code to compute static over-approximations of runtime behavior and represent them as program invariants over suitable domains.
+Figure 13: Incorrect invariant for bdd_manager class
+6 Related Work
+In this section, we discuss how ClassInvGen relates to previous works on synthesizing program invariants statically, dynamically, and neurally.
+6.1 Static approaches
+Static techniques, such as interpolation [31] or abstract interpretation [7] perform a symbolic analysis of source code to compute static over-approximations of runtime behavior and represent them as program invariants over suitable domains.
 These techniques are often used to prove the safety properties of the code.
-They focus on synthesizing loop invariants and method pre/postconditions, and a few around module-level specifications (Lahiri et al., 2009).
+They focus on synthesizing loop invariants and method pre/postconditions, and a few also handle module-level specifications [22].
 Given the undecidability of program verification, these techniques scale poorly for real-world programs, especially in the presence of complex data structures and frameworks.
 In contrast, ClassInvGen can be applied to large codebases to synthesize high-quality class invariants but does not guarantee soundness by construction.
-6.2. Dynamic approaches
-Dynamic synthesis techniques, such as Daikon (Ernst et al., 2007), DIG (Nguyen et al., 2012), SLING (Le et al., 2019), and specification mining (Ammons et al., 2002), learn invariants by observing the dynamic behaviors of programs over a set of concrete execution traces.
+6.2 Dynamic approaches
+Dynamic synthesis techniques, such as Daikon [11], DIG [36], SLING [25], and specification mining [1], learn invariants by observing the dynamic behaviors of programs over a set of concrete execution traces.
 One advantage of these dynamic techniques is that they can be agnostic to the code and generally applicable to different languages.
 However, these approaches are limited by the templates or patterns over which the invariants can be expressed.
-DySy (Csallner et al., 2008) employs dynamic symbolic execution to alleviate the problem of fixed templates for bounded executions but resorts to ad-hoc abstraction for loops or recursion.
-Hellendoorn et al.
-(2019) trained models to predict the quality of invariants generated by tools such as Daikon, but do not generate new invariants.
-SpecFuzzer (Molina et al., 2022) generates numerous candidate assertions via fuzzing to construct templates and filters them using Daikon and mutation testing.
-Finally, Geminus (Boockmann and Lüttgen, 2024) aims at synthesizing sound and complete class invariants representing the set of reachable states, guiding their search using random test cases termed Random Walk.
+DySy [8] employs dynamic symbolic execution to alleviate the problem of fixed templates for bounded executions but resorts to ad-hoc abstraction for loops or recursion.
+[17] trained models to predict the quality of invariants generated by tools such as Daikon, but they do not generate new invariants.
+SpecFuzzer [33] generates numerous candidate assertions via fuzzing to construct templates and filters them using Daikon and mutation testing.
+Finally, Geminus [2] aims at synthesizing sound and complete class invariants representing the set of reachable states, guiding their search using random test cases termed Random Walk.
 Unlike these approaches, ClassInvGen can generate a much larger class of invariants, leveraging multimodal inputs, including source code, test cases, comments, and even the naming convention learned from training data, to enhance invariant synthesis.
-Further, unlike prior dynamic approaches, LLM-based test generation (an active area of research (Lemieux et al., 2023; Schäfer et al., 2023; Yang et al., 2024)) reduces the need to have a high-quality test suite to obtain the invariants.
-For the use case of static verification, learning-based approaches have been used to iteratively improve the quality of the synthesized inductive invariants (Garg et al., 2014, 2016; Padhi et al., 2016) from dynamic traces.
+Further, unlike prior dynamic approaches, LLM-based test generation (an active area of research [27, 40, 45]) reduces the need to have a high-quality test suite to obtain the invariants.
+For the use case of static verification, learning-based approaches have been used to iteratively improve the quality of the synthesized inductive invariants [14, 15, 37] from dynamic traces.
 However, these approaches have not been evaluated in real-world programs due to the need for symbolic reasoning.
-6.3. Neural approaches
+6.3 Neural approaches
 LLM-based invariant synthesis is an emerging area of research with some noteworthy recent contributions.
-Pei et al.
-(2023) trained a model for zero-shot invariant synthesis, which incurs high training costs and lacks feedback-driven repair.
+[38] trained a model for zero-shot invariant synthesis, which incurs high training costs and lacks feedback-driven repair.
 Their approach uses Daikon-generated invariants as both training data and ground truth, which can lead to spurious invariants.
-Prior work on nl2postcond (Endres et al., 2024) prompts LLMs to generate pre and postcondition of Python and Java benchmarks, illustrating LLMs’ ability to generate high-quality specifications.
-However, they do not prune incorrect invariants and do not generate class invariants that ClassInvGen does.
-It is an interesting future work to combine this work with ClassInvGen to generate complete class-level specifications including pre and postconditions for the public methods of the class.
-For static verification, recent works include the use of LLM for intent-formalization from natural language (Lahiri, 2024), and inferring specifications and inductive program invariants (Kamath et al., 2023; Ma et al., 2024).
+Prior work on nl2postcond [10] prompts LLMs to generate pre- and postconditions of Python and Java benchmarks, illustrating LLMs’ ability to generate high-quality specifications.
+However, they do not prune incorrect invariants and do not generate class invariants.
+Combine this work with ClassInvGen would be an interesting direction for future work.
+Two very recent neuro–symbolic pipelines extend LLM prompting to other kinds of specifications.
+[44] combines GPT-4 with bounded-model checking to infer loop invariants: the LLM enumerates candidate predicates, a BMC oracle filters them, and the surviving predicates are re-assembled into provable invariants, yielding a 97 % success rate on 316 numeric-loop benchmarks.
+[42] (AutoSpec) weaves static slicing and an off-the-shelf program verifier with LLM generation to synthesise function-level contracts; AutoSpec verifies 79 % of heterogeneous benchmarks plus an X.509 parser case study.
+Both systems rely on static or SMT-based oracles and target scalar loops or procedure specifications, whereas ClassInvGen tackles pointer-rich class/object invariants in idiomatic C++ and validates them chiefly through dynamic test-suite execution plus mutation testing.
+The different oracle allows our approach to scale to data-structure code bases where precise SMT models are hard to obtain.
+For static verification, recent works include the use of LLM for intent formalization from natural language [21], and inferring specifications and inductive program invariants [19, 30].
 None of these techniques scale to real-world programs due to the need for complex symbolic reasoning.
-7. Limitation
+7 Limitations
+At present, ClassInvGen judges an invariant’s correctness with the same test suite that the LLM co-generates alongside that invariant.
+This design keeps the pipeline fully automated, but it also risks co-adaptation: the model can drift toward invariants that merely fit the behaviours exercised by its own tests, overstating their generality.
 ClassInvGen uses generated tests for invariant pruning, but the test suite may include spurious tests that can incorrectly prune valid invariants.
@@
 ClassInvGen partially addresses this issue through compositional generation, breaking down the code into manageable parts.
-Ongoing advancements in LLMs, as highlighted in recent work (Liu et al., 2024; Gao et al., 2023), are also expected to mitigate this limitation.
-For future work, we plan to integrate invariant generation with the generation of formal specifications for member functions, enabling LLM a more comprehensive understanding of program behavior.
+Ongoing advancements in LLMs, as highlighted in recent work [28, 13], are also expected to mitigate this limitation.
+For future work, we plan to integrate invariant generation with the generation of formal specifications for member functions, enabling LLMs to gain a more comprehensive understanding of program behavior.
 Additionally, we aim to evaluate ClassInvGen on larger and more complex systems beyond Z3, demonstrating its scalability to diverse codebases.
-8. Conclusion
+8 Conclusion
 In this paper, we describe an approach to leverage LLMs and a lightweight mixed static/dynamic approach to synthesize class invariants.
@@
 Our technique is currently limited by an automated way to integrate the generated tests into the build system of the underlying repo, and the need for developers to validate the invariants.
-We envision that integrating ClassInvGen with the continuous integration (CI) and pull requests (PR) can aid in scaling the approach to more developers.
+We envision that integrating ClassInvGen with continuous integration (CI) and pull requests (PR) can aid in scaling the approach to more developers.
 In future work, we also plan to investigate incorporating developer feedback to repair or strengthen generated invariants.
-Acknowledgments
-We would like to thank Nikolaj Bjørner for his encouragement and help in inspecting the several generated invariants for the Z3 source code and provide feedback about their utility.
-Appendix A Prompt
+Appendix 0.A Prompt
 1You are an expert in creating program invariants from code and natural language.
@@
 Do not add assert or any form of assertions.
-Figure 22. ClassInvGen Generation system prompt: instruction and task description.
+Figure 14: ClassInvGen Generation system prompt: instruction and task description. Lines 3-7 give the formal definition of a class invariant (must hold before and after every public method).
 15Input Format:
@@
 514. If you can decompose a single invariant into smaller ones, try to output multiple invariants.
-Figure 23. ClassInvGen Generation system prompt: input-output format.
+Figure 15: ClassInvGen Generation system prompt: input-output format.
 Figure 14 and Figure 15 show the system prompt used by ClassInvGen for invariant-test co-generation.
@@
 5‘‘‘
-Figure 24. ClassInvGen Generation user prompt template.
+Figure 16: ClassInvGen Generation user prompt template.
 1You are an expert in repairing program invariants from code and natural language.
@@
 414. If you can decompose a single invariant into smaller ones, try to output multiple invariants.
-Figure 25. ClassInvGen Refinement system prompt: instruction and task description.
+Figure 17: ClassInvGen Refinement system prompt: instruction and task description.
 1Please fix the failed invariants given the feedback, tests and the original source code.
@@
 21{tests}
-Figure 26. ClassInvGen Refinement user prompt template.
+Figure 18: ClassInvGen Refinement user prompt template.
+Appendix 0.B Daikon Invariants Frequency Tables
+Table 7: Invariants for avl_tree, 11 public methods.
+Invariant Count
+this->root has only one value 4
+this->root._M_t has only one value 4
+this->root._M_t.__uniq_ptr_impl<AvlTree::Node, std::default_delete<AvlTree::Node> >._M_t has only one value 4
+this[0] has only one value 3
+this->n one of { 3, 4 } 3
+this[0] != null 2
+this->root != null 2
+this->root._M_t != null 2
+this->root._M_t.__uniq_ptr_impl<AvlTree::Node, std::default_delete<AvlTree::Node> >._M_t != null 2
+this->n one of { 1, 2, 3 } 1
+this->n one of { 0, 3 } 1
+this->n >= 1 1
+this->n == 3 1
+t.root has only one value 1
+t.root._M_t has only one value 1
+t.root._M_t.__uniq_ptr_impl<AvlTree::Node, std::default_delete<AvlTree::Node> >._M_t has only one value 1
+t.n == 3 1
+this->n == return 1
+Table 8: Invariants for red_black_tree, 11 public methods.
+Invariant Count
+this[0] != null 3
+this->root != null 3
+this->root._M_t != null 3
+this->root._M_t.__uniq_ptr_impl<RedBlackTree::Node, std::default_delete<RedBlackTree::Node> >._M_t != null 3
+this->n one of { 3, 4, 6 } 3
+this->root has only one value 3
+this->root._M_t has only one value 3
+this->root._M_t.__uniq_ptr_impl<RedBlackTree::Node, std::default_delete<RedBlackTree::Node> >._M_t has only one value 3
+this[0] has only one value 2
+this->n >= 0 1
+(No intersection exists) 1
+t.root has only one value 1
+t.root._M_t has only one value 1
+t.root._M_t.__uniq_ptr_impl<RedBlackTree::Node, std::default_delete<RedBlackTree::Node> >._M_t has only one value 1
+t.n == 3 1
+No intersection 1
+Table 9: Invariants for linked_list, 8 public methods.
+Invariant Count
+this[0] has only one value 5
+this->head has only one value 5
+this->head._M_t has only one value 5
+this->head._M_t.__uniq_ptr_impl<LinkedList::Node, std::default_delete<LinkedList::Node> >._M_t has only one value 5
+this[0] != null 4
+this->head != null 4
+this->head._M_t != null 4
+this->head._M_t.__uniq_ptr_impl<LinkedList::Node, std::default_delete<LinkedList::Node> >._M_t != null 4
+this->tail[] elements != null 2
+this->tail[].next elements != null 2
+this->tail[].next._M_t elements != null 2
+this->n >= 0 2
+this->n == 0 1
+this->tail[].data elements one of { 1 } 1
+this->tail[].data one of { [1] } 1
+this->n one of { 1 } 1
+this->n one of { 1, 2 } 1
+this->tail[].data elements >= 1 1
+this->tail[].data elements <= this->n 1
+this->tail[].data elements one of { 1, 2, 4 } 1
+this->tail[].data one of { [1], [2], [4] } 1
+this->n one of { 2, 3 } 1
+this->tail[].data == [3] 1
+this->tail[].data elements == 3 1
+Table 10: Invariants for binary_search_tree, 11 public methods.
+Invariant Count
+this->root has only one value 4
+this->root._M_t has only one value 4
+this->root._M_t.__uniq_ptr_impl<BinarySearchTree::Node, std::default_delete<BinarySearchTree::Node> >._M_t has only one value 4
+this->n one of { 0, 2, 3 } 3
+this->n one of { 0, 3 } 3
+this[0] has only one value 3
+this[0] != null 2
+this->root != null 2
+this->root._M_t != null 2
+this->root._M_t.__uniq_ptr_impl<BinarySearchTree::Node, std::default_delete<BinarySearchTree::Node> >._M_t != null 2
+t.root has only one value 1
+t.root._M_t has only one value 1
+t.root._M_t.__uniq_ptr_impl<BinarySearchTree::Node, std::default_delete<BinarySearchTree::Node> >._M_t has only one value 1
+t.n == 3 1
+(this->n == return) == (return == orig(this->n)) 1
+Table 11: Invariants for heap, 7 public methods.
+Invariant Count
+this->comp._M_invoker has only one value 7
+this->comp._Function_base._M_manager has only one value 7
+this[0] has only one value 6
+this->data has only one value 6
+this->data._Vector_base<int, std::allocator<int> >._M_impl has only one value 6
+this->comp has only one value 6
+this->comp._Function_base._M_functor has only one value 6
+this->comp._Function_base._M_functor._M_unused has only one value 6
+this[0] != null 3
+this->data != null 3
+this->data._Vector_base<int, std::allocator<int> >._M_impl != null 3
+this->comp != null 3
+this->comp._M_invoker != null 3
+this->comp._Function_base._M_functor != null 3
+this->comp._Function_base._M_functor._M_unused != null 3
+this->comp._Function_base._M_manager != null 3
+this->data._Vector_base<int, std::allocator<int> >._M_impl._Vector_impl_data._M_start[] elements >= 1 2
+this->data._Vector_base<int, std::allocator<int> >._M_impl._Vector_impl_data._M_start != null 1
+this->data._Vector_base<int, std::allocator<int> >._M_impl._Vector_impl_data._M_finish != null 1
+this->data._Vector_base<int, std::allocator<int> >._M_impl._Vector_impl_data._M_end_of_storage != null 1
+Table 12: Invariants for hash_table, 7 public methods.
+Invariant Count
+::__digits == "000102…6979899" 3
+::__tag == "" 3
+this[0] has only one value 3
+this->hash_function has only one value 3
+this->hash_function._Function_base._M_functor has only one value 3
+this->hash_function._Function_base._M_functor._M_unused has only one value 3
+this->hash_function._Function_base._M_functor._M_pod_data == "" 3
+this->load_factor == 0.75 3
+this->table has only one value 3
+this->table._Vector_base<… >._M_impl has only one value 2
+this->table._Vector_base<… >._M_impl has only one value 1
+this->_num_elements one of { 0, 1 } 1
+this->_size == 10 1
+this->table._Vector_base<… > > > > >._M_impl._Vector_impl_data._M_end_of_storage 1
+key._M_dataplus has only one value 1
+key._M_dataplus._M_p one of { "key1", "key2" } 1
+key._M_string_length == 4 1
+this->_num_elements >= 0 1
+this->_size >= 0 1
+this->_num_elements <= this->_size 1
+Table 13: Invariants for vector, 11 public methods.
+Invariant Count
+this[0] has only one value 9
+this->capacity == 5 7
+this->data has only one value 5
+this->data[] elements one of { 1, 2 } 5
+this->n == 2 4
+this->data[] == [1, 2] 4
+this->n in this->data[] 4
+this->data[] == [1] 3
+this->n == 0 2
+this->n one of { 1 } 2
+this->n in return[] 1
+return[] == [1, 2] 1
+return[] elements one of { 1, 2 } 1
+this->capacity == 0 1
+this->data == null 1
+this->n == 1 1
+this->n == v.n 1
+this->capacity == v.capacity 1
+v.data has only one value 1
+v one of { 1, 2 } 1
+this->capacity one of { 5 } 1
+this->data[] sorted by < 1
+this->n <= this->capacity 1
+this->n < this->capacity 1
+this->n one of { 2, 5 } 1
+this->data[] elements == 1 1
+this->data[] one of { [1], [1, 2] } 1
+return one of { 1, 2 } 1
+Table 14: Invariants for queue, 7 public methods.
+Invariant Count
+this->data has only one value 5
+this->maxSize == 10 5
+this[0] has only one value 4
+this->data[] == [10, 20, 30] 2
+this->data[] elements one of { 10, 20, 30 } 2
+this->head one of { 0, 1 } 2
+this->tail == 3 2
+this->n one of { 2, 3 } 2
+this->maxSize in this->data[] 2
+this[0] != null 2
+this->data != null 2
+this->data[] elements >= 0 2
+this->data[] sorted by < 2
+this->head < this->maxSize 2
+this->tail < this->maxSize 2
+this->data[] elements one of { 1, 2 } 2
+this->data[] one of { [1], [1, 2] } 2
+this->tail one of { 0, 1, 2 } 2
+this->tail in this->data[] 2
+this->head - this->tail + this->n == 0 1
+this->tail one of { 2, 3, 100 } 1
+this->maxSize one of { 10, 160 } 1
+this->n < this->maxSize 1
+this->head one of { 0, 50 } 1
+this->tail >= 0 1
+this->n <= this->maxSize 1
+this->head <= this->tail 1
+this->head one of { 0, 2 } 1
+this->n one of { 0, 1 } 1
+this->head == other.head 1
+this->maxSize == other.maxSize 1
+this->data[] == [7, 14] 1
+this->data[] elements one of { 7, 14 } 1
+this->head == 0 1
+other.data has only one value 1
+other.tail == 2 1
+this->head one of { 0, 1, 2 } 1
+this->head - this->tail + return == 0 1
+return one of { 0, 1, 2 } 1
+Table 15: Invariants for stack, 6 public methods.
+Invariant Count
+this->maxSize one of { 10, 160, 1280 } 4
+this->data[] sorted by < 4
+this->data[] elements < this->maxSize 4
+this->n < this->maxSize 3
+this[0] != null 3
+this->data != null 3
+this[0] has only one value 2
+this->data[] elements >= 0 2
+this->n one of { 0, 1, 2 } 1
+this->maxSize == other.maxSize 1
+this has only one value 1
+this->data has only one value 1
+this->n == 2 1
+this->maxSize == 10 1
+other.data has only one value 1
+other.data[] == [1, 2] 1
+other.data[] elements one of { 1, 2 } 1
+this->n <= this->maxSize 1
+Appendix 0.C Additional Implementation Details
+This appendix provides additional details and examples for the implementation of ClassInvGen.
+0.C.1 Code Instrumentation and Invariant Examples
+Figure 19 shows how a public method is instrumented with invariant checks, and Figure 20 shows an example of an incorrect invariant.
+1bool AvlTree::empty() {
+2 check_invariant();
+3 auto ret = empty_original();
+4 check_invariant();
+5 return ret;
+6}
+7
+8bool AvlTree::empty_original() {
+9 return n == 0;
+10}
+(a) AvlTree instrumented with invariants
+1void AvlTree::check_invariant() {
+2 std::function<bool(const std::unique_ptr<Node>&)>
+3 is_balanced = [&](const std::unique_ptr<Node>& node) -> bool {
+4 if (!node) return true;
+5 int left_height = height(node->left);
+6 int right_height = height(node->right);
+7 if (std::abs(left_height - right_height) > 1)
+8 return false;
+9 return is_balanced(node->left) &&
+10 is_balanced(node->right);
+11 };
+12 assert(is_balanced(root));
+13}
+(b) Example of a correct AvlTree class invariant
+Figure 19: AvlTree instrumentation and invariant example
+1void AvlTree::check_invariant() {
+2 assert(height(root) == get_height(root));
+3}
+Figure 20: Example of an incorrect invariant of AvlTree because there is no get_height method
+0.C.2 Test Generation Examples
+Figure 21 shows an example of a generated test suite used for filtering invariants.
+1int main() {
+2 // Test Case 1: Basic insertions and traversals
+3 {
+4 AvlTree tree;
+5 tree.insert(10); tree.insert(20); tree.insert(5);
+6 tree.in_order_traversal();
+7 tree.pre_order_traversal();
+8 }
+9 // Test Case 2: Size, height, empty checks
+10 {
+11 AvlTree tree;
+12 tree.insert(10); tree.insert(20);
+13 tree.size(); tree.height(); tree.empty();
+14 }
+15 // ... more test cases ...
+16}
+Figure 21: A test suite generated for AvlTree
+0.C.3 Refinement Details
+ClassInvGen implements a feedback loop for refining failing invariants. Figure 22(a) shows an error message, while Figures 22(b) and 23 show the BST invariant before and after refinement.
+1avl_tree.cpp: In lambda function:
+2avl_tree.cpp:21:16: error: use of
+3’is_balanced’ before deduction of ’auto’
+(a) gcc compiler error messages
+1void AvlTree::check_invariant() {
+2  auto is_bst = [&](const std::unique_ptr<Node>& node,
+3 const T& min, const T& max) -> bool {
+4 if (!node) return true;
+5 if (node->data <= min $|$
+6 node->data >= max) return false;
+7 return is_bst(node->left, min, node->data) &&
+8 is_bst(node->right, node->data, max);
+9 };
+10 assert(is_bst(root, std::numeric_limits<T>::min(),
+11 std::numeric_limits<T>::max()));
+12}
+(b) BST property before refinement
+1void AvlTree::check_invariant() {
+2  std::function<bool(const std::unique_ptr<Node>&, const T&, const T&)>
+3 is_bst = [&](const std::unique_ptr<Node>& node, const T& min, const T& max) -> bool {
+4 if (!node) return true;
+5 if (node->data <= min $|$ node->data >= max) return false;
+6 return is_bst(node->left, min, node->data) &&
+7 is_bst(node->right, node->data, max);
+8 };
+9 assert(is_bst(root, std::numeric_limits<T>::min(), std::numeric_limits<T>::max()));
+10}
+Figure 23: BST property after refinement: auto is changed to explicit declarations